### v1

# Verifiable Natural Language to Linear Temporal Logic Translation: A Benchmark Dataset and Evaluation Suite

Abstract
Empirical evaluation of state-of-the-art natural-language (NL) to temporal-logic (TL) translation systems reveals near-perfect performance on existing benchmarks. However, current studies measure only the accuracy of the translation of NL logic into formal TL, ignoring a system’s capacity to ground atomic propositions into new scenarios or environments. This is a critical feature, necessary for the verification of resulting formulas in a concrete state space. Consequently, most NL-to-TL translation frameworks propose their own bespoke dataset in which the correct grounding is known a-priori, inflating performance metrics and neglecting the need for extensible, domain-general systems. In this paper, we introduce Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a unifying benchmark that measures verification and verifiability of automated NL-to-LTL translation. The dataset consists of three unique state spaces and thousands of diverse natural language specifications and corresponding formal specifications in temporal logic. Moreover, the benchmark contains sample traces to validate the temporal logic expressions. While the benchmark directly supports end-to-end evaluation, we observe that many frameworks decompose the process into i) lifting, ii) grounding, iii) translation, and iv) verification. The benchmark provides ground truths after each of these steps to enable researches to improve and evaluate different substeps of the overall problem. Using the benchmark, we evaluate several state-of-the-art NL-to-TL translation models and frameworks, including nl2spec, NL2TL, NL2LTL, sequence-to-sequence translation, and various prompting techniques. Our evaluation reveals that existing work is capable of reliably performing lifting and translation with high accuracy but struggle with grounding the translation in a state space, which stems from the lack of datasets to drive the research. Moreover, we provide traces that can be used for integrating verification into translation process, unlocking further directions for future research. To encourage methodologically sound advances in verifiable NL-to-LTL translation approaches, we release VLTL-Bench here: https://www.kaggle.com/datasets/dubascudes/vltl-bench.
1 Introduction
Formal verification is essential for the safe deployment of autonomous robots [24, 23], cyber–physical controllers [14], and safety-critical software systems [2]. Verification first begins with a specification that defines intent in precise temporal-logic (TL) [29, 3]. However, human stakeholders typically articulate intent in ambiguous natural language (NL) [26, 16, 15] and the conversion of this NL to TL is a challenging and time-consuming process that requires human experts [31, 4, 25]. Due to this complexity, automated translation of NL to TL has emerged as a core research problem [6, 33, 12, 28]. Recently, neural seq-to-seq models [11, 20, 13], grammar-constrained decoders [21, 9], and large language models (LLMs) [30, 6, 8, 7] have all demonstrated promising results on benchmark corpora, with reported string-exact accuracies often exceeding $90\%$. Despite these gains, we find existing evaluations provide an incomplete — and sometimes misleading — picture of true semantic competence. Specifically, we observe that high translation accuracies typically report the correctness of lifted (Section 2.1) TL predictions. While these expressions are syntactically correct, and indeed are equivalent to the expected lifted TL, there is little investigation into a framework’s ability to ground the lifted atomic propositions into a world state space. This oversight can be attributed to the widespread lack of an accompanying concrete world state space in existing corpora.
Figure 1: Overview of our dataset synthesis and evaluation pipline for NL to LTL translation. The scenario configuration is used to construct grounded APs as concrete action/target combinations in a given scenario. We generate unique NL representations from diverse references to the concrete variables. These free-form and concrete AP representations are inserted into our sentence, LTL, and trace templates to generate each corpora. Each component of our datasets can be used individually or combined to facilitate the holistic evaluation on NL-to-LTL translation frameworks.
Table 1: Comparison of existing LTL benchmarks and VLTL-Bench. Our datasets exhibit full coverage of logical/linguistic complexity, scenario inclusion, and verification support. $\times$ indicates total lack, $\sim$ indicates a feature could be estimated with processing, and $\checkmark$ indicates representation.
Dataset Complexity Inclusion
Linguistic Logical Lifting System Model Grounding Verification
CW (MacGlashan et al. [18]) $\times$ $\sim$ $\sim$ $\times$ $\times$
GLTL (Gopalan et al. [10]) $\times$ $\sim$ $\sim$ $\times$ $\times$
Navi (Wang et al. [27]) $\sim$ $\times$ $\times$ $\times$
Conformal (Wang et al. [28]) $\sim$ $\sim$ $\times$
VLTL-Bench Warehouse
VLTL-Bench Traffic Light
VLTL-Bench Search and Rescue
Current approaches for automated NL-to-TL translation are characterized by the use of multiple components, each with a distinct purpose in the end-to-end translation process. As noted earlier, we categorize these components into either lifting, grounding, or translation. Many frameworks employ different combinations of these components, such as nl2spec [7], which presupposes a pre-lifted and pre-grounded natural language specification as input. For example, an input "Globally, grant 0 and grant 1 do not hold at the same time until it is allowed," is assumed to be accompanied by a dictionary that maps the natural language sub-string "grant 1" to a known atomic proposition $g_{1}$. Therefore, this framework lacks the components for lifting and grounding, but includes a human auditing feature, which we omit from our analysis of automated translation frameworks. Meanwhile, another SOTA approach NL2TL [6] is divided into two processes. First, an LLM - in this case GPT3.5 [1] - is prompted to extract the atomic propositions (APs) from the NL input, creating “lifted NL" (NL without complex propositions). This new, less complex representation is then translated by a trained sequence-to-sequence model, such as T5 [22]. For example, an NL specification for a system: “You have an apple until you eat it.", will be converted into the lifted NL: “prop1 until prop2.". Here, prop1 is “You have an apple" and prop2 is “you eat it". This lifted NL is then translated to the lifted TL: “prop1 $\cup$ prop2." and the lifted propositions would be replaced.
These two frameworks are representative of the general research effort to produce better NL-to-TL translation frameworks, and yet no current frameworks have a holistic and faithful approach to translation. Therefore, we note the need for a benchmark that includes the data required to evaluate all components of the end-to-end translation pipeline - lifting, grounding, translation, and verification - in order to provide a complete picture of a framework’s true performance. We compare existing datasets and benchmarks on the basis of the linguistic and logical complexity of the end-to-end input and output, as well as the inclusion of features required to measure the performance of each individual component in Table 1.
We address these gaps with the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a benchmark that grounds linear temporal logic (LTL) in a concrete world state space while broadening linguistic and logical coverage. VLTL-Bench introduces three new datasets which contain both raw and lifted versions of NL specifications, an AP-to-Condition dictionary, lifted and grounded LTL, and examples of satisfying and unsatisfying traces. The provision of each of these elements enables us to evaluate the four abstract components present in emerging NL-to-TL translation frameworks. Furthermore, the inclusion of example traces allows us to further evaluate the semantic consistency of translation and grounding attempts, resulting in a new verification accuracy performance metric, which measures whether an output translation holds on an a given trace. This feature permits consistency checking against real formal traces- a heretofore omitted feature of NL-to-TL translation corpora which will improve future explorations of NL-to-TL translation. In plausible real-world applications, a system may be given some example traces with which to grade its own translation attempts. This provides the opportunity to improve translation attempts without human auditing. We believe that the ability to measure the performance of individual components of a given framework will clarify the weaknesses of existing frameworks and guide future work. We outline our dataset synthesis and translation evaluation pipelines in Figure 1.
In summary, our contributions are: (i) a single, extensible benchmark and data synthesis pipeline for the evaluation of all current NL-to-TL translation components; (ii) a first-of-its-kind evaluation of the verification potential of existing frameworks, using our example satisfying and unsatisfying traces; and (iii) an empirical study that exposes both new failure modes in current translation methods and the severe accuracy decline when grounding is required. By providing a rigorous and extensible yardstick, VLTL-Bench can catalyze methodologically sound advances in NL-to-TL translation and, more broadly, in grounding large language models for formal reasoning tasks.
2 Background and Related Work
2.1 Preliminaries
In this section, we formally define a number of key terms necessary to describe and evaluate natural language-temporal logic translation systems. In particular, we have found that several terms are often conflated or co-defined in prior work, often leading to confusion regarding framework abilities. In order to provide a cogent description of these systems, as well as a robust evaluation, we define these terms as follows:
  1. Scenario: Referred to in existing work as the “World", “Environment", or “Space". A set $S$ of conditions appearing on a trace.
  2. Condition: In model checking, a condition is a uniquely-named Boolean variable $c_{i}$.
  3. Atomic Proposition: $\pi\in\Phi$, where $\Phi$ is the set of propositional variables in an LTL expression. During LTL verification, $\pi_{i}$ is assigned a value by matching with a condition $c\in S$.
  4. Lifting: $\lambda$: NL $\rightarrow\Phi$, extracting substrings corresponding to APs from natural language.
  5. Grounding: $g(\pi)=c$, replacing an abstract AP in an LTL expression with a condition $c\in S$.
  6. Translation: $\tau$: NL $\rightarrow$ LTL, converting a natural language string into a formal LTL expression.
  7. Verification: Given a trace $\sigma$ or Kripke structure $K$, check whether a grounded LTL expression $g(\varphi)$ holds. For trace-based verification, construct a minimally satisfactory $K$ from $\sigma$.
2.2 Linear Temporal Logic
The syntax of LTL is given by the following grammar:
$\varphi::={}$ $\pi\;\mid\;\neg\varphi\;\mid\;\varphi_{1}\wedge\varphi_{2}\;\mid\;\varphi_{1}\vee\varphi_{2}\;\mid\;\varphi_{1}\Rightarrow\varphi_{2}$
$\;\mid\;\bigcirc\varphi\;\mid\;\diamondsuit\varphi\;\mid\;\Box\varphi\;\mid\;\varphi_{1}\,\cup\,\varphi_{2}$
We further discuss model checking with linear temporal logic in Appendix A.1 and Appendix A.2.
2.3 Benchmark Datasets
The Cleanup World (CW) dataset MacGlashan et al. [18] pairs NL commands with reward functions in an OO-MDP. While this work introduces a supervised grounding approach, the dataset used for training and evaluation is limited in terms of logical and linguistic complexity. This is due to the minimal set of APs used to construct the examples, and lack of diversity in both AP representation and sentence structure. Similarly, GLTL Gopalan et al. [10] gathers English sentences mapped to Geometric LTL formulas in robot-navigation scenarios, emphasizing stochastic-window operators but remaining modest in scale. The contents are remarkably similar to Cleanup World in both quality and metrics, with both datasets having the same number of unique temporal logic phrases and unique APs. The use of LLMs to enhance sentence diversity had not yet been adopted in the field at this time, and no alternate AP representations are introduced in the natural language. Therefore, the evaluation abilities of the GLTL dataset are limited in largely the same manner as Cleanup World. The Navi. corpus, introduced by Wang et al. [27] couples NL commands with LTL formulas and execution traces in a grid world. Its 226 unique APs make it a strong test of translation and lexical robustness, though this improvement comes at the cost of grounding- the Navi corpus does not specify formal APs, providing instead POS-tagged natural language representations, as shown below. Most recently, the ConformalNL2LTL dataset, introduced by Wang et al. [28], frames translation as open-vocabulary QA wrapped in conformal prediction (CP) to provide statistical guarantees. If the cumulative CP $p$-value falls below a user-specified threshold, the system escalates to a human; otherwise, it commits the subformula. The authors release a 1000-example corpus of grounded robot tasks partitioned by difficulty. We provide examples and a quantitative evaluation of existing datasets in Appendix A.3.
3 Verifiable Linear Temporal Logic Benchmark
Our framework is built around two core stages: Grounded Scenario Configuration, which defines the domain in which every proposition lives, and Data Synthesis, which instantiates our library of expert templates with scenario-specific atoms and then prunes any invalid NL. In the first stage, we provide the canonical formalization of the agent and the world it operates in. In the second stage, we apply the canonical APs obtained from the first to our expert-defined NL, TL, and trace templates.
3.1 Grounded Scenario Configuration
Each dataset is parameterized by a scenario—a small, declarative world model that provides:
  • Targets $\mathcal{L}$: e.g. {location, item} or {street, light}.
  • Actions $\mathcal{A}_{\mathrm{args}}$: verbs the agent may perform (e.g. idle, deliver(item, location), search(item)).
  • Target types $t\in\mathcal{T}$: each with its own action set $\mathcal{A}_{t}$ and argument signatures (e.g. deliver(item,location), avoid(threat)).
By swapping the scenario description of $\mathcal{L}$, $\mathcal{A}_{\mathrm{args}}$, and $\{\mathcal{A}_{t}\}$, the same logical and linguistic templates can be deployed in multiple domains without changing the core synthesis code.
3.2 Data Synthesis
To produce our datasets, we started with the 36 expert-crafted lifted NL-LTL pairs introduced by Cosler et al. [7] in the nl2spec benchmark and we crafted 7 of our own. We then crafted 43 pairs of corresponding example traces to each NL-LTL pair. Following this, we transform these lifted examples into a set of templates that can be used for diverse NL-LTL synthesis, introducing three substantial increases in complexity.
  1. 1.
    We generate more diverse words for objects, actions, and items as a potential scenario to enhance each example.
  2. 2.
    Each LTL skeleton (e.g. globally(prop_1 implies prop_2)) is coupled to action and target references from these scenarios.
  3. 3.
    Each example is given a corresponding pair of trace templates that yield a minimal satisfying ($\sigma_{good}\models\varphi_{G}$) and violating ($\,\sigma_{bad}\not\models\varphi_{G}$) trace.
To generate one example:
  1. 1.
    Atom sampling. Draw $k$ unique atomic propositions by randomly selecting actions and arguments from the scenario’s $\mathcal{A}_{\mathrm{ego}}\cup\mathcal{A}_{t}$ and $\mathcal{L}$.
  2. 2.
    Template instantiation. Uniformly choose a lifted template whose arity $\leq k$. Fill its LTL skeleton with the $k$ canonical atoms to obtain the grounded formula $\varphi_{g}$, and replace each atom by $\mathsf{prop}_{i}$ to obtain the lifted formula $\varphi_{L}$.
  3. 3.
    NL realization. Fill the semantic pattern with each atom’s surface-form (including articles/prepositions), apply morphological fixes (gerunds, capitalization), and record token-level spans. Emit both the free-form sentence and its grounded_sentence with explicit $\mathsf{prop}_{i}$ placeholders.
  4. 4.
    Trace filling. Apply the template’s trace builder to the list $[\mathsf{prop}_{1},\dots,\mathsf{prop}_{k}]$, yielding one good trace (satisfies $\varphi_{G}$) and one bad trace (violates $\varphi_{G}$).
  5. 5.
    Pruning. Submit each candidate NL sentence to GPT-4o for a lightweight check of the natural language components, removing any entries that the model flags as nonsensical or incorrect from the benchmark set.
3.3 Metrics
Each dataset entry includes three paired artifacts:
$\bigl{\{}\,\underbrace{\mathrm{sentence},\,\mathrm{lifted\ sentence\}}}_{\text{NL (grounded \& lifted)}},\;\underbrace{\varphi_{G},\,\varphi_{L}}_{\text{LTL(grounded \& lifted)}},\;\underbrace{\sigma_{good}\models\varphi_{G},\,\sigma_{bad}\not\models\varphi_{G}}_{\text{Traces (holds \& $\neg$ holds)}}\bigr{\}}.$
This rich annotation supports four independent evaluation axes, displayed in Figure 2:
  1. 1.
    Lifting accuracy: The ability to correctly extract substrings from a natural language specification that map to a corresponding atomic proposition in the linear temporal logic.
  2. 2.
    Translation accuracy: The fidelity of mapping grounded or lifted NL to the correct temporal-logic formula.
  3. 3.
    Grounding accuracy: The ability to align logical atomic proposition labels with their corresponding world state condition.
  4. 4.
    Verification accuracy: The capacity to decide satisfaction or violation of a grounded LTL formula on a given trace or traces.
3.4 Datasets
To use our data synthesis pipeline, we provide three scenario definitions (listed in Appendix A.5) accompanied by action and target references. In total, our synthesis pipeline yields three new datasets, usable for training and evaluation. Though the same generation pipeline is used to produce all three datasets, we note that the input and In this section, we describe each of the included datasets.
Figure 2: Overview of an isolated evaluation of each individual component. Lifting accuracy measures accuracy of predicted natural language AP spans, grounding accuracy measures the performance on mapping AP spans to world state conditions, translation accuracy measures the performance on NL-LTL translation on the token-level, and verification accuracy is an approach to measuring whether a grounded LTL expression holds on a trace.
Warehouse.
Our Warehouse dataset simulates a realistic warehouse retrieval scenario, explicitly designed for scalability and complexity in grounding tasks. Warehouse is our most distinct dataset with its inclusion of all 80 COCO [17] object classes, significantly enriching the semantic and linguistic complexity and variation of atomic propositions (APs). Each entry includes LTL formulas with explicit grounding and alignment at token-level granularity. The dataset also uniquely provides verified positive ("good") and negative ("bad") execution traces for robust validation.
Example:
  • Sentence: “At every moment, at least one of drop off the long chair to the loading dock, wait, or look for the glass for alcoholic beverage holds.”
  • Lifted Sentence: “At every moment, at least one of prop_1, prop_2 or prop_3 holds.”
  • Grounded LTL Formula: globally( deliver(bench, loading_dock) or idle or search(wine_glass))
  • APs: prop_1 = “drop off long chair to loading dock”, prop_2 = “idle”, prop_3 = “look for glass for alcoholic beverage”
Search and Rescue.
The Search-and-Rescue (S&R) scenario presents a detailed context of emergency response and robotic aid. It emphasizes multi-step TL relationships, requiring precise event sequencing and robust conditional propositions. Like the Warehouse scenario, it provides complete grounding information and includes scenario-specific LTL templates with both positive and negative execution traces for rigorous verification. The S&R scenario deliberately includes an "avoid" action for the purpose of stress-testing the a framework’s ability to distinguish between APs and LTL operators, because avoid is commonly used to indicate a "globally (negation $\varphi$)" pattern.
Example:
  • Sentence: “If deliver aid to the safe rescuer holds and request assistance holds next, then stay away from the active gas leak holds in the step after that.”
  • Lifted Sentence: “If prop_1 holds and prop_2 holds next, then prop_3 holds in the step after that.”
  • Grounded LTL Formula: globally(deliver_aid(safe_rescuer) and next get_help implies next next avoid(active_gas_leak))
  • APs: prop_1 = deliver aid to safe rescuer, prop_2 = request assistance, prop_3 = avoid active gas leak
Traffic Light Control.
The Traffic Light Control scenario introduces LTL propositions relevant to urban navigation and infrastructure management. This scenario tests an NL-to-LTL system’s ability to accurately ground complex real-world relationships, such as spatial landmarks and persistent behaviors like "infinitely often." Each dataset entry is provided with explicit good and bad traces, detailed proposition grounding, and is constructed via a flexible scenario template, allowing straightforward extension to additional real-world tasks.
Example:
  • Sentence: “Take a photo of the motorcycle on northeast 10th avenue must keep holding until, from some point on, change east light red holds infinitely often.”
  • Lifted Sentence: “prop_1 must keep holding until prop_2 holds infinitely often.”
  • Grounded LTL Formula: photo(motorcycle,northeast_10th_avenue) until globally(finally change(light_east,red))
  • APs: prop_1 = take a photo of motorcycle, northeast 10th avenue, prop_2 = change east light red
4 Experimental Results
In this section, we present the results of multiple evaluations of NL-to-LTL translation frameworks and components. In Section 4.1, we measure the performance of common natural language lifting approaches, evaluated on four existing datasets in addition to the three datasets we present in VLTL-Bench. In Section 4.2, we evaluate three SOTA NL-to-LTL frameworks on lifted NL to lifted TL translation. Note here, that measuring lifted translation performance on existing datasets is particularly difficult, as they present varying degrees of clarity in their lifted natural language elements. The CW [18], GLTL [10], and Navi [27] datasets have been processed to include lifted natural language components by Chen et al. [6] - and we perform similar processing of the Conformal dataset [28] to include it in our evaluation. In Section 4.3 we develop and evaluate two grounding baselines on our three datasets. In Section 4.4, we assemble the best results from the three individual evaluations to perform the first end-to-end translation evaluation. In Section 4.5 we perform our novel verification evaluation over the example traces of our dataset.
4.1 Lifting Evaluation
First, we evaluate four language models on the natural language lifting task. The LLM-based approaches each use the lifting prompt template from the NL2TL framework [6], which includes few-shot ground-truth NL to lifted NL examples from each of the datasets. The input to both models is a natural-language sentence and we compare the prediction made by the model against the ground-truth lifted natural language and measure the average IoU scores over every individual NL AP. We present the results in Table 2 where we see the linguistic complexity of our datasets is highlighted in the accuracies as even the best scoring model (GPT-4.1) reduces in performance on our new datasets. While this performance drop is minor on this model, it is more significant on the lower-cost, smaller GPT models. Regardless, this indicates our success in increasing the evaluation complexity.
Table 2: Comparison of lifting approaches. This table displays the mean IoU between predicted lifted natural language APs and ground-truth lifted natural language APs.
Accuracy (%)
S&R Traffic Light Warehouse
Model GLTL [10] CW [18] Conformal [28] Navi [27] (ours) (ours) (ours)
GPT-3.5-turbo 81.6 78.6 76.8 71.0 65.3 59.4 67.9
GPT-4o-mini 84.9 82.3 85.6 81.0 66.7 63.1 68.9
GPT-4.1-mini 97.7 95.9 96.1 97.1 94.4 96.6 93.1
4.2 Lifted Translation Evaluation
Next, we evaluate the lifted translation capabilities of the three NL-to-LTL frameworks - nl2spec, NL2LTL, and NL2TL. In order to analyze the performance of their lifted translation abilities, the ground-truth lifted NL specification is given to the translation model, and the resulting lifted LTL translation is compared against the ground-truth lifted LTL. We perform this evaluation as a binary semantic accuracy using the pyModelChecking library [5] and present the results in Table 3. Here, we see that lifted translation can be very successful with both out-of-the-box LLM prompting (nl2spec) and with fine-tuned seq2seq models. However, as we have noted, we will see in end-to-end evaluation that this is an overconfident estimation of translation performance as grounding is not considered.
Table 3: Binary Accuracy on the Lifted NL to Lifted LTL task across seven datasets, for three frameworks (NL2LTL, nl2spec, NL2TL) given ground-truth lifted NL specifications.
Accuracy (%)
S&R Traffic Light Warehouse
Framework Translation Model GLTL [10] CW [18] Conformal [28] Navi [27] (ours) (ours) (ours)
GPT-3.5-turbo 37.9 48.1 18.3 9.9 11.9 13.2 13.8
NL2LTL [8] GPT-4o-mini 38.6 55.4 23.6 10.4 12.3 13.9 12.5
GPT-4.1-mini 51.7 64.6 42.1 39.7 41.6 40.0 37.4
GPT-3.5-turbo 44.4 40.9 35.2 50.3 51.1 46.3 50.2
nl2spec [7] GPT-4o-mini 77.3 80.1 73.5 69.7 74.9 75.8 74.2
GPT-4.1-mini 89.8 92.9 78.3 81.5 89.1 91.6 88.4
NL2TL [6] t5-base 99.9 99.9 94.9 99.7 100.0 100.0 100.0
4.3 Grounding Evaluation
In this section, we present the results obtained from our evaluation of our baseline grounding framework, applied to the ground truth lifted TL from our three VLTL-Bench datasets. We use two prompting strategies (described in Appendix A.4) applied to three GPT models to provide a broad evaluation of current grounding capabilities. Our first prompting baseline - few-shot - is composed of a brief description of the task at hand, accompanied by nine few-shot examples of correct (sentence, lifted sentence, AP-dictionary) tuples from all three scenarios (as opposed to individual scenarios). The next strategy is the scenario baseline prompt which includes the full scenario configuration file, as well as three few-shot examples from the dataset. To measure grounding accuracy, we parse the resulting AP-dictionary predictions and compare them with our ground-truth knowledge of the AP-dictionary in each entry. Our metrics are per-AP and per-AP-dictionary accuracy. Per-AP accuracy is calculated by recording the total number of correctly grounded APs divided by the total number of APs in the test set, and per-AP-dictionary accuracy is calculated by recording the total number of correct complete AP-dictionaries, divided by the size of the test set. These results are presented in Table 4.
Our evaluation of the two grounding baselines reveals that even advanced LLMs struggle to accurately ground lifted APs into a concrete world state space - even when the parameters of this state space are provided, as is done in the scenario baseline. We observe that even though the scenario baseline achieves lower performance on most benchmarks and settings, it beats the few-shot baseline on our Warehouse scenario when comparing the more powerful reasoning models. As noted in Section 3, the Warehouse scenario is specifically designed to stress-test grounding and lifting. We conclude that the provision of the world state space in the scenario baseline includes information that aids reasoning models in determining which world state conditions are referred to in the lifted APs, but the overall performance of these baselines on the grounding task remains notably lower than other tasks involved in verifiable NL-to-LTL translation.
Table 4: Comparison of Grounding approaches. This table displays binary accuracy between predicted AP Grounding and known AP dictionary. LLM Baseline uses 9 few–shot sentence + lifted sentence + AP dict examples from every dataset; “Scenario” includes the scenario definition in the prompt and 3 examples from only that dataset.
Accuracy (of APs) Accuracy (of AP Dictionaries)
Prompt Model S&R Traffic Light Warehouse S&R Traffic Light Warehouse
GPT-3.5-turbo 56.9 69.5 18.3 34.2 51.4 7.4
Few-shot GPT-4o-mini 82.3 66.5 18.4 68.6 48.4 7.0
GPT-4.1-mini 77.3 67.4 23.8 60.4 45.8 7.8
GPT-3.5-turbo 76.7 37.3 13.6 63.6 20.8 5.0
Scenario GPT-4o-mini 66.7 44.8 23.6 44.8 16.8 9.2
GPT-4.1-mini 68.6 27.9 34.4 45.2 15.4 13.0
4.4 End-to-End Translation Evaluation
Now, we perform and end-to-end evaluation which considers the accumulation of the three individual translation steps. For all three frameworks, we select the best-performing component (model) from each of the individual evaluations (lifting, grounding, and translation) to assemble an end-to-end translation framework which factors in the combined performance of all the translation steps. We see in Table 5, that as a result of the poor grounding results of all current approaches, the high performance of the lifting and lifted translation steps is diminished, resulting in a poor overall semantic accuracy of the final translation. Our datasets show that even the best performing model (NL2TL) does not approach real-world performance needs, inciting the need for NL-to-TL translation approaches which consider a concrete world state space.
Table 5: End-to-end evaluation of all three SOTA frameworks using the best lifting, translation, and grounding components. We report the binary accuracy of the resulting LTL.
Accuracy (%)
Framework S&R Traffic Light Warehouse
NL2LTL [8] 35.4 38.4 26.2
nl2spec [7] 34.8 33.6 29.6
NL2TL [6] 54.4 60.1 46.2
4.5 Verification Evaluation
Finally, we present the results of our experiments on the verification of LTL outputs from each of the three NL-to-LTL translation frameworks that we compare. We use the outputs from our lifted translation evaluation (Table 3) to isolate the verification metric from the lifting task, and apply our LLM-baseline grounding frameworks. In Table 6, out results demonstrate that even frameworks exhibiting accurate lifted NL to lifted TL translation suffer a notable decline in performance when grounding relies on systems similar to our LLM baselines. Furthermore, this evaluation supports the use of trace satisfaction in place of ground-truth LTL comparison as a metric for grounded translation accuracy, because the example traces encode the minimum specifications of correctly grounded and translated LTL. In future frameworks, example traces could be used as part of a feedback loop to grounding and translation components.
Table 6: Performance (binary accuracy) on S&R, Traffic Light, and Warehouse, broken down into satisfied holding traces, satisfied not-holding traces, and both. All three frameworks are evaluated on both grounding strategies using their top-scoring lifted translation model.
S&R Traffic Light Warehouse
Framework Grounding Strategy Sat Unsat Both Sat Unsat Both Sat Unsat Both
NL2LTL [8] Few-shot 61.6 61.4 35.4 64.6 60.2 38.4 52.4 58.6 26.2
Scenario 1.06 32.0 7.4 61.8 59.2 36.6 12.4 36.2 9.8
nl2spec [7] Few-shot 47.4 48.0 34.8 47.2 46.0 33.6 46.0 44.2 29.6
Scenario 34.0 36.4 21.0 40.2 41.8 28.2 32.0 34.6 19.0
NL2TL [6] Few-shot 75.0 79.4 54.4 80.2 80.6 60.8 71.4 74.8 46.2
Scenario 27.5 50.8 22.1 72.6 76.3 54.5 33.3 52.4 23.5
5 Conclusion
We present the Verifiable Linear Temporal Logic Benchmark. VLTL-Bench is a suite of three new NL-to-LTL translation datasets that include the standard natural language and LTL pairs, supplemented with lifted natural language, lifted LTL, and trace examples. These additional features provide a method for the isolated training and evaluation of individual NL-to-LTL translation framework components. The provision of trace examples in VLTL-Bench introduces the possibility of a new type of input that is plausible in real-world translation frameworks, but unrepresented in current corpora. We acknowledge that the datasets included in the VLTL-Bench suite are generated using a finite number of linguistic and logical templates, populated by diverse synthetic natural language APs. Therefore, our logical diversity is limited to 43 lifted NL-LTL templates. The linguistic complexity of our benchmark is also partially limited in the same manner. However, our randomly selected action and target references combined with LLM-based pruning allows this work to surpass existing datasets in terms of diversity and realism, and we believe future effort should be partially devoted to further improvements in this area. VLTL-Bench reveals significant weaknesses in what were previously ironclad NL-to-LTL translation frameworks. Among these our the reliance on accurately lifted NL inputs for translation, lack of accurate grounding components, and lack of example trace inputs in current approaches. We hope that contribution will encourage exploration of diverse methods for NL-to-LTL translation, beyond the use of LLMs.
Appendix A Appendix
In this appendix, we present a detailed overview of linear temporal logic in A.1, a discussion of verification via Kripke structures in A.2, a quantitative comparison of our VLTL-Bench dataset against existing datasets as well as examples from those datasets in A.3, our developed prompts for the baseline grounding approaches in A.4, the configuration files for our three scenarios in A.5, and finally our estimated compute resource usage and our external code and license information in A.6.
A.1 Linear Temporal Logic
Linear temporal logic (LTL) is a modal extension of classical propositional logic that enables reasoning about how truths evolve over a discrete, linear timeline [32]. Formulas in LTL are interpreted over infinite sequences (or “traces”) of states
$\sigma=s_{0},s_{1},s_{2},\ldots,$
where each state $s_{i}$ (which has a set of conditions) specifies which atomic propositions $\pi^{\mu}$ hold true at time $i$. This framework makes it possible to specify and verify both safety properties (e.g., “nothing bad ever happens”) and liveness properties (e.g., “something good eventually happens”), and it underpins many model-checking techniques for reactive systems.
The syntax of LTL is given by the following grammar:
$\varphi::={}$ $\pi\;\mid\;\neg\varphi\;\mid\;\varphi_{1}\wedge\varphi_{2}\;\mid\;\varphi_{1}\vee\varphi_{2}\;\mid\;\varphi_{1}\Rightarrow\varphi_{2}$
$\;\mid\;\bigcirc\varphi\;\mid\;\diamondsuit\varphi\;\mid\;\Box\varphi\;\mid\;\varphi_{1}\,\cup\,\varphi_{2}$
[…]
$\sigma,i\models\varphi_{1}\cup\varphi_{2}$ $\quad\text{iff}\quad\exists k\geq i:\;\sigma,k\models\varphi_{2}\;\wedge\;\forall j\in[i,k):\;\sigma,j\models\varphi_{1}.$
Although our focus is on discrete-time LTL, many of these ideas carry over to related formalisms such as signal temporal logic (STL) for continuous-time, real-valued signals [19].
A.2 Verification via Kripke Structures and Fluents
Verification of LTL specifications is typically conducted using a Kripke structure, which is a formal transition system comprising states, transitions, and labels indicating which atomic propositions hold true in each state. Formally, a Kripke structure is defined as a tuple $M=(S,S\_0,R,L)$, where:
  • $S$ is a finite set of states,
  • $S_{0}\subseteq S$ is the set of initial states,
  • $R\subseteq S\times S$ is the transition relation, specifying allowed state transitions,
[…]
Verification involves checking whether every possible path through the Kripke structure satisfies the given LTL formula. For instance, safety properties such as “a collision never occurs” require that no path through the structure contains a state labeled with the proposition collision. Conversely, liveness properties such as “a goal is eventually reached” demand the existence of a future state in every valid path labeled with the proposition goal. Additionally, verification explicitly involves fluents—timestamped state variables that indicate when certain conditions or states become true. Each fluent captures both the state variable (atomic proposition) and the time step at which the transition into the corresponding state occurs. Formally, a fluent can be represented as a tuple $(\pi^{\mu},t)$, indicating that proposition $\pi^{\mu}$ becomes true at time step $t$ due to a state transition within the Kripke structure. Fluents bridge the gap between high-level temporal specifications and lower-level state transitions, facilitating practical model checking and control synthesis in robot control systems.
A.3 Existing Datasets
Table 7: Comparison of NL–LTL datasets. Note that VLTL-Bench datasets include satisfying and unsatisfying traces for each entry. VLTL-Bench separates the total data into training and testing, as well as a subset of data taken from training and testing partitions that was selected by GPT-4o for its particularly realistic natural language.
Dataset Size Unique TL # APs # Actions # Args
GLTL Gopalan et al. [10] 11,109 37 4 1 4
CW MacGlashan et al. [18] 3,371 37 4 1 4
Conformal Wang et al. [28] 1,000 212 239 4 235
Navi Wang et al. [27] 7,474 6,414 221 26
Search-and-rescue [VLTL-Bench] 7,304 5,425 220 7 44
Traffic-light [VLTL-Bench] 7,319 6,196 5,046 4 175
Warehouse [VLTL-Bench] 7,457 5,991 5,074 5 452
Cleanup World (CW).
  • Sentence: “go to the blue room keep going and stop when you reach the green room”
  • LTL Formula: “finally(blue_room and finally green_room)”
  • Grounded Sentence: “go to the prop_1 keep going and stop when you reach the green prop_2,”
  • APs: prop_1 = go to blue room, prop_2 = go to green room.
GLTL.
  • Sentence: “enter the blue or red room and proceed until the green room”
  • LTL Formula: “finally((red_room or blue_room) and finally green_room)”
  • Grounded Sentence: “enter the prop_2 or prop_1 and proceed until the green prop_3,”
  • APs: prop_1 = go to red room, prop_2 = go to blue room, prop_3 = go to green room
Navi.
  • Sentence: “at some time get hold apple or whenever acquire pear”
  • LTL Formula: “finally(get_hold_v apple_n or finally(acquire_v pear_n)”
  • Grounded Sentence: “at some time prop_1 or whenever prop_2”
  • APs: prop_1 = get_hold_v apple_n, prop_2 = acquire_v pear_n
ConformalNL2LTL.
  • Sentence: “Stay in parking lot 4 until you reach car 5”
  • LTL Formula: “parking_lot_4 until car_5”
  • Grounded Sentence: “Stay in prop_1 until you reach prop_2”
  • APs: prop_1 = go to parking lot 4, prop_2 = go to car 5
A.4 Grounding Prompts
This section includes the few-shot examples used in our grounding prompt baselines. The few-shot baselines uses all of the following in its prompt, while the scenario baseline includes only the scenario specific few-shot examples combined with the scenario description, given in Appendix A.5
Figure 3: Few-shot examples for the Warehouse Scenario
Figure 4: Few-shot examples for the Traffic Light Scenario
Figure 5: Few-shot examples for the Search and Rescue Scenario
A.5 Scenario Configurations
In this section, we provide the scenario configuration files that are inserted into the grounding prompts and used for data generation.
Figure 6: Warehouse Scenario Configuration file
Figure 7: Traffic Light Scenario Configuration file
Figure 8: Search and Rescue Scenario Configuration file
A.6 Compute Resources and External Code and License Information
All LLM inference was performed using the OpenAI API. Approximately \$30.00 in compute credits were used for our evaluations. The T5-base model used by NL2TL was trained and tested locally on a machine using an Nvidia GeForce RTX 4070Ti Super 16 GB GPU, an Intel i9 14900KF, and 64 GB of RAM. Training took approximately 40 minutes using a batch size of 16 and a learning rate of $2\text{e}^{-5}$ for 3 epochs.
The nl2spec framework is released at https://github.com/realChrisHahn2/nl2spec under the MIT license, the NL2TL framework is released at https://github.com/yongchao98/NL2TL?tab=readme-ov-file with no attached license, the NL2LTL framework is released at https://github.com/IBM/nl2ltl under the MIT license, and the pyModelChecking library is released at https://github.com/albertocasagrande/pyModelChecking under the GNU General Public License.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All claims in the beginning of the paper are justified by the experiments and details provided in the main body of the paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Limitations are discussed in Section 5.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: We do not provide any theories. All results are experimental, and all experiments are provided.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: The paper is thorough in its description of the experiments and the data. Someone familiar with the area could feasibly replicate this work given the citations to existing work and the details provided. Regardless, we provide all code and the dataset is public and provided as required.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: The dataset is public and the code and experimental details will be in the supplemental material.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Full details are in the code, but the information in the paper is sufficient for understanding the work and the meaning of the evaluations.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: Our evaluations present only the mean of each metric. Our evaluations are (relatively) small, composed of <1000 examples, and we focus on the presentation and descriptions of our new datasets, rather than a detailed review of current frameworks.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: See Appendix A.6.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We have no ethical issues and have read the code.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss the benefits to NL to TL translation which can be used for improved safe operation of formally specified systems.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A] .
  44. Justification: Our dataset does not enable dangerous models or model behavior.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: See Appendix A.6.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes] .
  52. Justification: Dataset is published as required. All code released in the supplemental material is well documented.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: No human subjects
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: No human subjects
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [Yes]
  64. Justification: All uses of LLM for help with dataset generation and for experimental evaluation of methods which use LLMs were disclosed. In addition we have provided the prompts for our unique usage.

### v2

@@
 Abstract
-Empirical evaluation of state-of-the-art natural-language (NL) to temporal-logic (TL) translation systems reveals near-perfect performance on existing benchmarks.
-However, current studies measure only the accuracy of the translation of NL logic into formal TL, ignoring a system’s capacity to ground atomic propositions into new scenarios or environments.
+Empirical evaluation of state-of-the-art natural language (NL) to temporal logic (TL) translation systems reveals near-perfect performance on existing benchmarks.
+However, current studies only measure the accuracy of the translation of NL logic into formal TL, ignoring a system’s capacity to ground atomic propositions into new scenarios or environments.
 This is a critical feature, necessary for the verification of resulting formulas in a concrete state space.
-Consequently, most NL-to-TL translation frameworks propose their own bespoke dataset in which the correct grounding is known a-priori, inflating performance metrics and neglecting the need for extensible, domain-general systems.
-In this paper, we introduce Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a unifying benchmark that measures verification and verifiability of automated NL-to-LTL translation.
-The dataset consists of three unique state spaces and thousands of diverse natural language specifications and corresponding formal specifications in temporal logic.
-Moreover, the benchmark contains sample traces to validate the temporal logic expressions.
+In this paper, we introduce the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a unifying benchmark for automated NL-to-LTL translation.
+The dataset consists of three unique state spaces and thousands of diverse natural language specifications and their corresponding formal temporal logic specifications.
+Moreover, the benchmark contains sample traces to verify the temporal logic expressions.
 While the benchmark directly supports end-to-end evaluation, we observe that many frameworks decompose the process into i) lifting, ii) grounding, iii) translation, and iv) verification.
-The benchmark provides ground truths after each of these steps to enable researches to improve and evaluate different substeps of the overall problem.
-Using the benchmark, we evaluate several state-of-the-art NL-to-TL translation models and frameworks, including nl2spec, NL2TL, NL2LTL, sequence-to-sequence translation, and various prompting techniques.
-Our evaluation reveals that existing work is capable of reliably performing lifting and translation with high accuracy but struggle with grounding the translation in a state space, which stems from the lack of datasets to drive the research.
-Moreover, we provide traces that can be used for integrating verification into translation process, unlocking further directions for future research.
-To encourage methodologically sound advances in verifiable NL-to-LTL translation approaches, we release VLTL-Bench here: https://www.kaggle.com/datasets/dubascudes/vltl-bench.
+The benchmark provides ground truths after each of these steps to enable researchers to improve and evaluate different substeps of the overall problem.
+Using the benchmark, we evaluate several state‑of‑the‑art NL-to-TL translation models and frameworks, including nl2spec, NL2TL, NL2LTL, Lang2LTL, sequence-to-sequence translation, and various LLM prompting techniques.
+Our evaluation confirms that existing work is capable of reliably performing lifting and translation with high accuracy, while it exposes their struggles to ground the translation into a state space, which stems from the lack of existing datasets.
 1 Introduction
-Formal verification is essential for the safe deployment of autonomous robots [24, 23], cyber–physical controllers [14], and safety-critical software systems [2].
-Verification first begins with a specification that defines intent in precise temporal-logic (TL) [29, 3].
-However, human stakeholders typically articulate intent in ambiguous natural language (NL) [26, 16, 15] and the conversion of this NL to TL is a challenging and time-consuming process that requires human experts [31, 4, 25].
-Due to this complexity, automated translation of NL to TL has emerged as a core research problem [6, 33, 12, 28].
-Recently, neural seq-to-seq models [11, 20, 13], grammar-constrained decoders [21, 9], and large language models (LLMs) [30, 6, 8, 7] have all demonstrated promising results on benchmark corpora, with reported string-exact accuracies often exceeding $90\%$.
-Despite these gains, we find existing evaluations provide an incomplete — and sometimes misleading — picture of true semantic competence.
-Specifically, we observe that high translation accuracies typically report the correctness of lifted (Section 2.1) TL predictions.
-While these expressions are syntactically correct, and indeed are equivalent to the expected lifted TL, there is little investigation into a framework’s ability to ground the lifted atomic propositions into a world state space.
-This oversight can be attributed to the widespread lack of an accompanying concrete world state space in existing corpora.
-Figure 1: Overview of our dataset synthesis and evaluation pipline for NL to LTL translation.
-The scenario configuration is used to construct grounded APs as concrete action/target combinations in a given scenario.
-We generate unique NL representations from diverse references to the concrete variables.
-These free-form and concrete AP representations are inserted into our sentence, LTL, and trace templates to generate each corpora.
-Each component of our datasets can be used individually or combined to facilitate the holistic evaluation on NL-to-LTL translation frameworks.
-Table 1: Comparison of existing LTL benchmarks and VLTL-Bench.
-Our datasets exhibit full coverage of logical/linguistic complexity, scenario inclusion, and verification support.
-$\times$ indicates total lack, $\sim$ indicates a feature could be estimated with processing, and $\checkmark$ indicates representation.
-Dataset Complexity Inclusion
-Linguistic Logical Lifting System Model Grounding Verification
-CW (MacGlashan et al. [18]) $\times$ $\sim$ $\sim$ $\times$ $\times$
-GLTL (Gopalan et al. [10]) $\times$ $\sim$ $\sim$ $\times$ $\times$
-Navi (Wang et al. [27]) $\sim$ $\times$ $\times$ $\times$
-Conformal (Wang et al. [28]) $\sim$ $\sim$ $\times$
-VLTL-Bench Warehouse
-VLTL-Bench Traffic Light
-VLTL-Bench Search and Rescue
-Current approaches for automated NL-to-TL translation are characterized by the use of multiple components, each with a distinct purpose in the end-to-end translation process.
-As noted earlier, we categorize these components into either lifting, grounding, or translation.
-Many frameworks employ different combinations of these components, such as nl2spec [7], which presupposes a pre-lifted and pre-grounded natural language specification as input.
-For example, an input "Globally, grant 0 and grant 1 do not hold at the same time until it is allowed," is assumed to be accompanied by a dictionary that maps the natural language sub-string "grant 1" to a known atomic proposition $g_{1}$.
-Therefore, this framework lacks the components for lifting and grounding, but includes a human auditing feature, which we omit from our analysis of automated translation frameworks.
-Meanwhile, another SOTA approach NL2TL [6] is divided into two processes.
-First, an LLM - in this case GPT3.5 [1] - is prompted to extract the atomic propositions (APs) from the NL input, creating “lifted NL" (NL without complex propositions).
-This new, less complex representation is then translated by a trained sequence-to-sequence model, such as T5 [22].
-For example, an NL specification for a system: “You have an apple until you eat it.", will be converted into the lifted NL: “prop1 until prop2.".
-Here, prop1 is “You have an apple" and prop2 is “you eat it".
-This lifted NL is then translated to the lifted TL: “prop1 $\cup$ prop2." and the lifted propositions would be replaced.
-These two frameworks are representative of the general research effort to produce better NL-to-TL translation frameworks, and yet no current frameworks have a holistic and faithful approach to translation.
-Therefore, we note the need for a benchmark that includes the data required to evaluate all components of the end-to-end translation pipeline - lifting, grounding, translation, and verification - in order to provide a complete picture of a framework’s true performance.
-We compare existing datasets and benchmarks on the basis of the linguistic and logical complexity of the end-to-end input and output, as well as the inclusion of features required to measure the performance of each individual component in Table 1.
-We address these gaps with the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a benchmark that grounds linear temporal logic (LTL) in a concrete world state space while broadening linguistic and logical coverage.
-VLTL-Bench introduces three new datasets which contain both raw and lifted versions of NL specifications, an AP-to-Condition dictionary, lifted and grounded LTL, and examples of satisfying and unsatisfying traces.
-The provision of each of these elements enables us to evaluate the four abstract components present in emerging NL-to-TL translation frameworks.
-Furthermore, the inclusion of example traces allows us to further evaluate the semantic consistency of translation and grounding attempts, resulting in a new verification accuracy performance metric, which measures whether an output translation holds on an a given trace.
-This feature permits consistency checking against real formal traces- a heretofore omitted feature of NL-to-TL translation corpora which will improve future explorations of NL-to-TL translation.
-In plausible real-world applications, a system may be given some example traces with which to grade its own translation attempts.
-This provides the opportunity to improve translation attempts without human auditing.
-We believe that the ability to measure the performance of individual components of a given framework will clarify the weaknesses of existing frameworks and guide future work.
-We outline our dataset synthesis and translation evaluation pipelines in Figure 1.
-In summary, our contributions are: (i) a single, extensible benchmark and data synthesis pipeline for the evaluation of all current NL-to-TL translation components; (ii) a first-of-its-kind evaluation of the verification potential of existing frameworks, using our example satisfying and unsatisfying traces; and (iii) an empirical study that exposes both new failure modes in current translation methods and the severe accuracy decline when grounding is required.
-By providing a rigorous and extensible yardstick, VLTL-Bench can catalyze methodologically sound advances in NL-to-TL translation and, more broadly, in grounding large language models for formal reasoning tasks.
+Formal verification is essential for the safe deployment of autonomous robots (RobotLanguage; Spec2), cyber-physical controllers (SurveyTemporal), and safety-critical software systems (SAS_textbook_PoCPS).
+Verification first begins with a specification that defines intent in precise temporal logic (TL) (watson2005autonomous; bellini2000temporal).
+However, human stakeholders typically articulate intent in ambiguous natural language (NL) (veizaga2021systematically; lamar2009linguistic; lafi2021eliciting), and the conversion of this NL to TL is a challenging and time-consuming process that requires human experts (yin2024formal; cardoso2021review; thistle1986control).
+Due to this complexity, automated NL-to-TL translation has emerged as a core research problem (NL2TL; Natural2CTL_data; circuits_data; conformalLTL_data).
+Recently, neural sequence-to-sequence models (hahn2022formal; pan2023data; hsiung2022generalizing), grammar-constrained decoders (post2018fastlexicallyconstraineddecoding; geng2024grammarconstraineddecodingstructurednlp), and large language models (LLMs) (xu2024learning; NL2TL; NL2LTL; nl2spec) have all demonstrated promising results on benchmark corpora, with reported accuracies often exceeding $90\%$.
+Despite these gains, evaluations are misleading as most datasets only test lifted translation, where temporal logic formulas contain abstract placeholders for atomic propositions (APs).
+The harder task of grounded translation—instantiating APs with domain-specific actions and arguments—is usually left unmeasured.
+This imbalance stems from limitations of current datasets, which omit the annotations required to separately evaluate lifting, translation, and grounding.
+As a result, current frameworks optimize for partial tasks, leaving open the more difficult but necessary problem of grounding for producing fully executable specifications.
+Figure 1: Overview of our dataset synthesis and evaluation framework for NL-to-LTL translation.
+The framework used a configuration file to define concrete and unique scenarios.
+The data synthesis generates the NL and TL pairs with associated traces for verification while providing ground truth results for intermediate components.
+Benchmarks for NL-to-TL translation include CW (CW_data), GLTL (GLTL-data), Navi (Navi_data), and Conformal (conformalLTL_data).
+Their limitations are fourfold.
+(i) Although recent frameworks decompose the task into lifting, translation, and grounding, these benchmarks supply ground truth only for the end-to-end result (NL-TL pairs), preventing assessment of intermediate components.
+(ii) CW and GLTL omit grounding entirely, yielding translations without executable semantics.
+For example, the NL specification: “Go to the green room and then go to the blue room.” is mapped to the LTL expression: $\diamondsuit G\rightarrow\diamondsuit B$”, without providing a grounded definition of the predicates $G$ and $B$.
+(iii) Navi and Conformal nominally support grounding but rely on overly simplistic state spaces (e.g., Navi’s colored-room grid), which fails to capture the referential and contextual ambiguities of natural language.
+(iv) Execution traces/trajectories for independent semantic verification (e.g., via model checking), are not provided, preventing rigorous evaluation.
+In this paper, we introduce the Verifiable Linear Temporal Logic Benchmark (VLTL-Bench), a benchmark that grounds linear temporal logic (LTL) in a concrete world state space while broadening linguistic and logical coverage through more diverse atomic propositions.
+As illustrated in Figure 1, VLTL-Bench exposes every stage of the NL-to-TL pipeline: raw and lifted NL specifications, an AP-to-Reference dictionary, lifted and grounded LTL formulas, and both satisfying and unsatisfying traces.
+Our dataset synthesis and evaluation framework for NL-to-LTL translation leverages scenario configurations to construct grounded action/target combinations, from which we synthesize diverse natural language representations and integrate them into sentence, LTL, and trace templates, yielding corpora whose components can be used individually or combined for holistic evaluation.
+This layered design makes it possible to isolate performance on lifting, translation, grounding, and verification individually, while also enabling end-to-end evaluation.
+We provide four scenario configuration files and construct a Kitchen Assistant, Traffic light, Search & Rescue, and Warehouse dataset.
+Using these four datasets we evaluate the capabilities and limitations of state-of-the-art NL to TL translation frameworks.
+In summary, we propose: (i) a single, extensible benchmark for evaluating all NL-to-TL translation components; (ii) the first verification evaluation using satisfying and unsatisfying traces; and (iii) an empirical study that reveals both new failure modes in current methods and the severe accuracy decline when grounding is required.
+The remainder of this paper is organized as follows.
+Section 2 covers preliminaries for LTL and model checking.
+Section 3 contains a detailed description of the Verifiable Linear Temporal Logic Benchmark datasets, as well as details on how they were synthesized.
+Section 4 includes an evaluation of current NL-to-TL frameworks on both Verifiable Linear Temporal Logic Benchmark and existing datasets.
+We conclude our paper in 5.
+Additional details may be found in the Appendix A
 2 Background and Related Work
+In this section we introduce necessary notation and background information on temporal logic systems including terminology, linear temporal logic symbols, and existing NL-to-TL datasets.
+Linear Temporal Logic. The syntax of LTL is given by the following grammar:
+$\varphi::={}$ $\pi\;\mid\;\neg\varphi\;\mid\;\varphi_{1}\wedge\varphi_{2}\;\mid\;\varphi_{1}\vee\varphi_{2}\;\mid\;\varphi_{1}\Rightarrow\varphi_{2}$
+$\;\mid\;\bigcirc\varphi\;\mid\;\diamondsuit\varphi\;\mid\;\Box\varphi\;\mid\;\varphi_{1}\,\cup\,\varphi_{2}$
+We further discuss model checking with linear temporal logic in Appendix A.1 and Appendix A.2.
 2.1 Preliminaries
-In this section, we formally define a number of key terms necessary to describe and evaluate natural language-temporal logic translation systems.
-In particular, we have found that several terms are often conflated or co-defined in prior work, often leading to confusion regarding framework abilities.
+In this section, we formally define a number of key terms necessary to describe and evaluate NL-to-TL translation systems.
 In order to provide a cogent description of these systems, as well as a robust evaluation, we define these terms as follows:
-  1. Scenario: Referred to in existing work as the “World", “Environment", or “Space". A set $S$ of conditions appearing on a trace.
+  1. Scenario: Referred to in existing work as the “World”, “Environment”, or “Space”. A set $S$ of conditions appearing on a trace.
   2. Condition: In model checking, a condition is a uniquely-named Boolean variable $c_{i}$.
@@
 For trace-based verification, construct a minimally satisfactory $K$ from $\sigma$.
-2.2 Linear Temporal Logic
-The syntax of LTL is given by the following grammar:
-$\varphi::={}$ $\pi\;\mid\;\neg\varphi\;\mid\;\varphi_{1}\wedge\varphi_{2}\;\mid\;\varphi_{1}\vee\varphi_{2}\;\mid\;\varphi_{1}\Rightarrow\varphi_{2}$
-$\;\mid\;\bigcirc\varphi\;\mid\;\diamondsuit\varphi\;\mid\;\Box\varphi\;\mid\;\varphi_{1}\,\cup\,\varphi_{2}$
-We further discuss model checking with linear temporal logic in Appendix A.1 and Appendix A.2.
-2.3 Benchmark Datasets
-The Cleanup World (CW) dataset MacGlashan et al.
-[18] pairs NL commands with reward functions in an OO-MDP.
-While this work introduces a supervised grounding approach, the dataset used for training and evaluation is limited in terms of logical and linguistic complexity.
-This is due to the minimal set of APs used to construct the examples, and lack of diversity in both AP representation and sentence structure.
-Similarly, GLTL Gopalan et al.
-[10] gathers English sentences mapped to Geometric LTL formulas in robot-navigation scenarios, emphasizing stochastic-window operators but remaining modest in scale.
-The contents are remarkably similar to Cleanup World in both quality and metrics, with both datasets having the same number of unique temporal logic phrases and unique APs.
-The use of LLMs to enhance sentence diversity had not yet been adopted in the field at this time, and no alternate AP representations are introduced in the natural language.
-Therefore, the evaluation abilities of the GLTL dataset are limited in largely the same manner as Cleanup World.
+Table 1: Comparison of existing LTL benchmarks and VLTL-Bench.
+We report the number of unique words across all NL specifications and the number of unique LTL specifications.
+Additionally, we report support for lifting, grounding, and verification.
+Dataset # Words # TL Lifting Translation Grounding Verification
+CW (CW_data) 184 37 $\sim$ $\times$ $\times$
+GLTL (GLTL-data) 183 37 $\sim$ $\times$ $\times$
+Navi (Navi_data) 131 6414 $\times$ $\sim$ $\times$
+Conformal (conformalLTL_data) 439 212 $\sim$ $\checkmark$ $\sim$ $\times$
+VLTL-Bench Warehouse 1028 5991
+VLTL-Bench Traffic Light 217 6196
+VLTL-Bench Search and Rescue 245 5425
+VLTL-Bench Kitchen Assistant 376 7385
+2.2 Existing Benchmark Datasets
+In this section, we review existing benchmarks for NL-to-TL translation.
+We compare these corpora in terms of linguistic and logical complexity, and support for evaluation of different framework modules in Table 2.1.
+We measure the complexity using the number of unique words appearing in natural language specifications (#Words), as well as the number of unique temporal logic expressions (#TL).
+In terms of modules, we report if a dataset has support for evaluation of lifting, grounding, and verification.
+We also provide examples from existing datasets in Appendix A.5.
+In Table 2.1, we observe that the older datasets Cleanup World (CW) (CW_data) and GLTL (GLTL-data) from the pre-LLM era have limited complexity both in terms of unique words and temporal logic expressions.
+While they support evaluation of translation, the lifting data is not explicitly given, the APs do not vary in their form to any meaningful degree, and they can be lexically identified in both the NL and TL elements of each entry ("green room" $\leftrightarrow G$, "blue room" $\leftrightarrow B$, etc.).
 The Navi.
-corpus, introduced by Wang et al.
-[27] couples NL commands with LTL formulas and execution traces in a grid world.
-Its 226 unique APs make it a strong test of translation and lexical robustness, though this improvement comes at the cost of grounding- the Navi corpus does not specify formal APs, providing instead POS-tagged natural language representations, as shown below.
-Most recently, the ConformalNL2LTL dataset, introduced by Wang et al.
-[28], frames translation as open-vocabulary QA wrapped in conformal prediction (CP) to provide statistical guarantees.
-If the cumulative CP $p$-value falls below a user-specified threshold, the system escalates to a human; otherwise, it commits the subformula.
-The authors release a 1000-example corpus of grounded robot tasks partitioned by difficulty.
-We provide examples and a quantitative evaluation of existing datasets in Appendix A.3.
-3 Verifiable Linear Temporal Logic Benchmark
-Our framework is built around two core stages: Grounded Scenario Configuration, which defines the domain in which every proposition lives, and Data Synthesis, which instantiates our library of expert templates with scenario-specific atoms and then prunes any invalid NL.
-In the first stage, we provide the canonical formalization of the agent and the world it operates in.
-In the second stage, we apply the canonical APs obtained from the first to our expert-defined NL, TL, and trace templates.
+corpus, introduced by (Navi_data), couples NL commands with LTL formulas in a grid world.
+As Table 2.1 shows, Navi exhibits a substantial increase in logical complexity, with 6,414 unique formulas and support for partial grounding.
+Its 221 unique APs make it a strong test of translation and lexical robustness, though this improvement comes at the cost of well-defined grounding rules: the corpus does not specify formal APs, providing instead POS-tagged natural language representations.
+As reflected in Table 2.1, the Conformal (conformalLTL_data) dataset introduces 439 unique words and 212 formulas with explicit grounding, but its scale is modest at 1,000 examples.
+In contrast, VLTL-Bench provides a testbed suited to holistic evaluation across lifting, translation, grounding, and verification.
+We provide a more detailed quantitative comparison between these datasets and VLTL-Bench in Section 3.4.
+3 The Verifiable Linear Temporal Logic Benchmark
+In the following subsections, we first introduce Grounded Scenario Configuration, which formalizes the world model by defining types, targets, and actions that ensure well-typed logical atoms.
+We then describe our Data Synthesis pipeline, which instantiates expert-crafted NL–LTL templates with scenario-specific atoms to produce paired sentences, formulas, and traces.
+Next, we present the Metrics used to evaluate each stage of the NL-to-LTL pipeline, and finally, we detail the Datasets generated from three scenario definitions, highlighting their unique challenges and properties.
 3.1 Grounded Scenario Configuration
+To formalize how natural language specifications map onto executable logical structures, we distinguish three interconnected components: types, targets, and actions.
+Types serve as abstract categories that describe what kinds of objects or entities an action can take as input (e.g., a location, an item, or a threat).
+Targets are the grounded instantiations of these action–type combinations, where abstract slots are filled with concrete constants.
+Actions are verbs that capture the capabilities of the agent; each action comes with a signature that specifies the expected types of its arguments.
+Together, this hierarchy ensures that linguistic expressions can be systematically mapped into well-typed logical atoms: types constrain argument structure, actions define the permissible predicates, and targets bind them to domain-specific instances.
 Each dataset is parameterized by a scenario—a small, declarative world model that provides:
-  • Targets $\mathcal{L}$: e.g. {location, item} or {street, light}.
-  • Actions $\mathcal{A}_{\mathrm{args}}$: verbs the agent may perform (e.g. idle, deliver(item, location), search(item)).
-  • Target types $t\in\mathcal{T}$: each with its own action set $\mathcal{A}_{t}$ and argument signatures (e.g. deliver(item,location), avoid(threat)).
-By swapping the scenario description of $\mathcal{L}$, $\mathcal{A}_{\mathrm{args}}$, and $\{\mathcal{A}_{t}\}$, the same logical and linguistic templates can be deployed in multiple domains without changing the core synthesis code.
+Types $t\in\mathcal{T}$: denotes the sort of parameters accepted by an action (e.g. item or location).
+Targets $\mathcal{L}$: Specific instances of typed arguments, (e.g. an argument apple of type item, or an argument loading_dock of type location).
+Actions $\mathcal{A}_{\mathrm{args}}$: verbs the agent may perform, which may have one or more targets, (e.g. idle() has no targets, deliver(apple, loading_dock) takes two—item and location).
 3.2 Data Synthesis
-To produce our datasets, we started with the 36 expert-crafted lifted NL-LTL pairs introduced by Cosler et al.
-[7] in the nl2spec benchmark and we crafted 7 of our own.
-We then crafted 43 pairs of corresponding example traces to each NL-LTL pair.
-Following this, we transform these lifted examples into a set of templates that can be used for diverse NL-LTL synthesis, introducing three substantial increases in complexity.
+To produce our datasets, we began with the 36 expert-crafted lifted NL-LTL pairs of the nl2spec benchmark (nl2spec), and we added 7 new ones of our own (provided in Appendix A.4).
+We then transformed these 43 examples into templates to support diverse NL–LTL synthesis.
+Finally, for each NL–LTL example, we crafted one pair of traces—one satisfying and one violating.
+Each dataset entry includes a tuple of these three artifacts,
+[…]
+and is algorithmically constructed with the following steps:
   1. 1.
-    We generate more diverse words for objects, actions, and items as a potential scenario to enhance each example.
+    Template selection. Uniformly choose a lifted template. Each template has an arity that determines how many atomic propositions must be instantiated.
   2. 2.
-    Each LTL skeleton (e.g. globally(prop_1 implies prop_2)) is coupled to action and target references from these scenarios.
-  3. 3.
-    Each example is given a corresponding pair of trace templates that yield a minimal satisfying ($\sigma_{good}\models\varphi_{G}$) and violating ($\,\sigma_{bad}\not\models\varphi_{G}$) trace.
-To generate one example:
-  1. 1.
-    Atom sampling. Draw $k$ unique atomic propositions by randomly selecting actions and arguments from the scenario’s $\mathcal{A}_{\mathrm{ego}}\cup\mathcal{A}_{t}$ and $\mathcal{L}$.
-  2. 2.
-    Template instantiation.
-Uniformly choose a lifted template whose arity $\leq k$.
-Fill its LTL skeleton with the $k$ canonical atoms to obtain the grounded formula $\varphi_{g}$, and replace each atom by $\mathsf{prop}_{i}$ to obtain the lifted formula $\varphi_{L}$.
+    Atom sampling.
+For each argument slot in the template, draw a unique atomic proposition by randomly selecting actions and arguments from the scenario’s $\mathcal{A}_{t}$ and $\mathcal{L}$.
+Let $k$ denote the total number of sampled atoms.
+Fill the LTL skeleton with these $k$ atoms to obtain the grounded formula $\varphi_{G}$, and replace each atom by $\mathsf{prop}_{i}$ to obtain the lifted formula $\varphi_{L}$.
   3. 3.
     NL realization.
-Fill the semantic pattern with each atom’s surface-form (including articles/prepositions), apply morphological fixes (gerunds, capitalization), and record token-level spans.
-Emit both the free-form sentence and its grounded_sentence with explicit $\mathsf{prop}_{i}$ placeholders.
+Fill the template pattern with each atom’s surface form (including articles/prepositions), apply morphological fixes (gerunds, capitalization), and record token‐level spans.
+Emit both the free‐form sentence and its grounded_sentence with explicit $\mathsf{prop}_{i}$ placeholders.
   4. 4.
     Trace filling.
-Apply the template’s trace builder to the list $[\mathsf{prop}_{1},\dots,\mathsf{prop}_{k}]$, yielding one good trace (satisfies $\varphi_{G}$) and one bad trace (violates $\varphi_{G}$).
-  5. 5.
-    Pruning.
-Submit each candidate NL sentence to GPT-4o for a lightweight check of the natural language components, removing any entries that the model flags as nonsensical or incorrect from the benchmark set.
+Apply the template’s trace patterns to the list $\mathsf{prop}_{1},\dots,\mathsf{prop}_{k}]$, yielding one positive trace (satisfies $\varphi_{G}$) and one negative trace (violates $\varphi_{G}$).
+This rich annotation supports four independent evaluation axes, displayed in Figure 2.
 3.3 Metrics
-Each dataset entry includes three paired artifacts:
-$\bigl{\{}\,\underbrace{\mathrm{sentence},\,\mathrm{lifted\ sentence\}}}_{\text{NL (grounded \& lifted)}},\;\underbrace{\varphi_{G},\,\varphi_{L}}_{\text{LTL(grounded \& lifted)}},\;\underbrace{\sigma_{good}\models\varphi_{G},\,\sigma_{bad}\not\models\varphi_{G}}_{\text{Traces (holds \& $\neg$ holds)}}\bigr{\}}.$
-This rich annotation supports four independent evaluation axes, displayed in Figure 2:
-  1. 1.
-    Lifting accuracy: The ability to correctly extract substrings from a natural language specification that map to a corresponding atomic proposition in the linear temporal logic.
-  2. 2.
-    Translation accuracy: The fidelity of mapping grounded or lifted NL to the correct temporal-logic formula.
-  3. 3.
-    Grounding accuracy: The ability to align logical atomic proposition labels with their corresponding world state condition.
-  4. 4.
-    Verification accuracy: The capacity to decide satisfaction or violation of a grounded LTL formula on a given trace or traces.
-3.4 Datasets
-To use our data synthesis pipeline, we provide three scenario definitions (listed in Appendix A.5) accompanied by action and target references.
-In total, our synthesis pipeline yields three new datasets, usable for training and evaluation.
-Though the same generation pipeline is used to produce all three datasets, we note that the input and In this section, we describe each of the included datasets.
+In this section, we introduce four complementary metrics that capture performance at different levels of the NL-to-TL pipeline, which is illustrated in Figure 2.
+Lifting accuracy measures the identification of atomic proposition spans in natural language, grounding accuracy evaluates their mapping to world state conditions, translation accuracy assesses logical equivalence between predicted and reference formulas, and verification accuracy checks whether predicted formulas satisfy or violate traces as expected.
+Together, these metrics provide a comprehensive view of system performance.
 Figure 2: Overview of an isolated evaluation of each individual component.
 Lifting accuracy measures accuracy of predicted natural language AP spans, grounding accuracy measures the performance on mapping AP spans to world state conditions, translation accuracy measures the performance on NL-LTL translation on the token-level, and verification accuracy is an approach to measuring whether a grounded LTL expression holds on a trace.
+Lifting accuracy.
+For each token $\mathbb{S}_{i}$ in a sentence, the system predicts a label $\hat{\lambda}(\mathbb{S}_{i})\in\{0,1,\dots,k\}$, where 0 denotes background and $n$ denotes membership in $\pi_{n}$.
+$\text{LiftAcc}\;=\;\frac{1}{|\mathbb{S}|}\sum_{i=1}^{|\mathbb{S}|}\bigl[\hat{\lambda}(\mathbb{S}_{i})=\lambda(\mathbb{S}_{i})\bigr].$
+This measures the token-level classification accuracy of mapping substrings to atomic propositions.
+Translation accuracy.
+Given a natural language specification $s$, the system produces a predicted TL formula $\hat{\varphi}$.
+Translation accuracy is an exact match between the predicted and reference formulas:
+$\text{TransAcc}\;=\;\bigl[\hat{\varphi}\equiv\varphi\bigr],$
+where $\equiv$ denotes logical equivalence. When working with lifted NL, the target is $\varphi_{L}$; for grounded NL, the target is $\varphi_{G}$.
+Grounding accuracy. Let $\{\texttt{prop}_{1},\dots,\texttt{prop}_{k}\}$ be lifted placeholders and $g_{\mathcal{S}}$ the gold grounding function. The system predicts $\hat{g}_{\mathcal{S}}$.
+$\text{GroundAcc}\;=\;\frac{1}{k}\sum_{j=1}^{k}\bigl[\hat{g}_{\mathcal{S}}(\texttt{prop}_{j})=g_{\mathcal{S}}(\texttt{prop}_{j})\bigr].$
+This measures how well predicted atoms match their reference predicates and arguments.
+Verification accuracy.
+For each dataset entry, two traces are provided: a positive trace $\sigma_{good}$ (satisfies $\varphi_{G}$) and a negative trace $\sigma_{bad}$ (violates $\varphi_{G}$).
+Given a predicted grounded formula $\hat{\varphi}_{G}$, verification checks whether the satisfaction relation holds:
+$\text{VerifAcc}\;=\;\frac{1}{2}\Big([\sigma_{good}\models\hat{\varphi}_{G}]\;+\;[\sigma_{bad}\not\models\hat{\varphi}_{G}]\Big).$
+3.4 Datasets
+We construct four scenario definitions accompanied by action and target references, namely a Kitchen Assistant, Traffic Light, Search & Rescue, and Warehouse scenario.
+The details are provided in Appendix A.8.
+Using our proposed data synthesis, we generate four new datasets for training and evaluation.
+Each of our four datasets is designed to highlight distinct challenges for NL-to-LTL translation: the Traffic Light Control scenario is intended to balance action and argument grounding challenges, including a large library of “street name” arguments, but a smaller set of actions; the Search-and-Rescue scenario emphasizes multi-step temporal dependencies and deliberately includes ambiguous actions such as “avoid” and “communicate” to stress-test the system’s ability to distinguish between natural language verbs and temporal operators; the Kitchen Assistant scenario includes intentionally ambiguous action and argument references in order to stress both action and argument grounding simultaneously; and the Warehouse scenario introduces high semantic and linguistic variability by incorporating all 80 COCO object classes, making grounding especially complex.
+In this section, we use an entry from the Warehouse dataset as an example to illustrate the structure and properties of our data; additional examples from the other scenarios are provided in Appendix A.7.
+Table 2: Comparison of NL–LTL datasets.
+We report the total number of entries (Size), the total number of unique TL entries, and the total number of unique APs appearing in the TL entries.
+Note that these datasets do not explicitly provide quantities of actions and arguments, and these are estimated by the authors.
+Dataset Size Unique TL # APs # Actions # Args
+GLTL GLTL-data 11,109 37 4 1 4
+CW CW_data 3,371 37 4 1 4
+Conformal conformalLTL_data 1,000 212 239 4 235
+Navi Navi_data 7,474 6,414 221 26
+Kitchen Assistant [VLTL-Bench] 10,000 7,385 172 9 36
+Search-and-rescue [VLTL-Bench] 7,304 5,425 220 7 44
+Traffic-light [VLTL-Bench] 7,319 6,196 5,046 4 175
+Warehouse [VLTL-Bench] 7,457 5,991 5,074 5 82
 Warehouse.
 Our Warehouse dataset simulates a realistic warehouse retrieval scenario, explicitly designed for scalability and complexity in grounding tasks.
-Warehouse is our most distinct dataset with its inclusion of all 80 COCO [17] object classes, significantly enriching the semantic and linguistic complexity and variation of atomic propositions (APs).
-Each entry includes LTL formulas with explicit grounding and alignment at token-level granularity.
-The dataset also uniquely provides verified positive ("good") and negative ("bad") execution traces for robust validation.
+Warehouse is our most distinct dataset with its inclusion of all 80 COCO (lin2014microsoft) object classes, significantly enriching the semantic and linguistic complexity and variation of atomic propositions.
+As with each of our datasets, all entries include LTL formulas with explicit grounding and alignment at token-level granularity, as well as verified positive (”good”) and negative (”bad”) execution traces for robust validation.
 Example:
@@
   • Lifted Sentence: “At every moment, at least one of prop_1, prop_2 or prop_3 holds.”
-  • Grounded LTL Formula: globally( deliver(bench, loading_dock) or idle or search(wine_glass))
-  • APs: prop_1 = “drop off long chair to loading dock”, prop_2 = “idle”, prop_3 = “look for glass for alcoholic beverage”
-Search and Rescue.
-The Search-and-Rescue (S&R) scenario presents a detailed context of emergency response and robotic aid.
-It emphasizes multi-step TL relationships, requiring precise event sequencing and robust conditional propositions.
-Like the Warehouse scenario, it provides complete grounding information and includes scenario-specific LTL templates with both positive and negative execution traces for rigorous verification.
-The S&R scenario deliberately includes an "avoid" action for the purpose of stress-testing the a framework’s ability to distinguish between APs and LTL operators, because avoid is commonly used to indicate a "globally (negation $\varphi$)" pattern.
-Example:
-  • Sentence: “If deliver aid to the safe rescuer holds and request assistance holds next, then stay away from the active gas leak holds in the step after that.”
-  • Lifted Sentence: “If prop_1 holds and prop_2 holds next, then prop_3 holds in the step after that.”
-  • Grounded LTL Formula: globally(deliver_aid(safe_rescuer) and next get_help implies next next avoid(active_gas_leak))
-  • APs: prop_1 = deliver aid to safe rescuer, prop_2 = request assistance, prop_3 = avoid active gas leak
-Traffic Light Control.
-The Traffic Light Control scenario introduces LTL propositions relevant to urban navigation and infrastructure management.
-This scenario tests an NL-to-LTL system’s ability to accurately ground complex real-world relationships, such as spatial landmarks and persistent behaviors like "infinitely often." Each dataset entry is provided with explicit good and bad traces, detailed proposition grounding, and is constructed via a flexible scenario template, allowing straightforward extension to additional real-world tasks.
-Example:
-  • Sentence: “Take a photo of the motorcycle on northeast 10th avenue must keep holding until, from some point on, change east light red holds infinitely often.”
-  • Lifted Sentence: “prop_1 must keep holding until prop_2 holds infinitely often.”
-  • Grounded LTL Formula: photo(motorcycle,northeast_10th_avenue) until globally(finally change(light_east,red))
-  • APs: prop_1 = take a photo of motorcycle, northeast 10th avenue, prop_2 = change east light red
+  • Grounded LTL Formula: globally( deliver(bench, loading_dock) or idle() or search(wine_glass))
+  • APs: prop_1 = “drop off long chair to loading dock”, prop_2 = “wait”, prop_3 = “look for glass for alcoholic beverage”
+  • Positive Trace: [deliver(bench, loading_dock)], [idle()], [search(wine_glass)]
+  • Negative Trace: [idle()], [idle()], [search(wine_glass), deliver(bench, loading_dock)]
 4 Experimental Results
 In this section, we present the results of multiple evaluations of NL-to-LTL translation frameworks and components.
-In Section 4.1, we measure the performance of common natural language lifting approaches, evaluated on four existing datasets in addition to the three datasets we present in VLTL-Bench.
+In Section 4.1, we measure the performance of common natural language lifting approaches, evaluated on four existing datasets in addition to three of the four datasets we present in VLTL-Bench.
 In Section 4.2, we evaluate three SOTA NL-to-LTL frameworks on lifted NL to lifted TL translation.
 Note here, that measuring lifted translation performance on existing datasets is particularly difficult, as they present varying degrees of clarity in their lifted natural language elements.
-The CW [18], GLTL [10], and Navi [27] datasets have been processed to include lifted natural language components by Chen et al.
-[6] - and we perform similar processing of the Conformal dataset [28] to include it in our evaluation.
+In both translation evaluations, we use the pyModelChecking library (pymodelcheck_code) to determine logical equivalence.
+The CW (CW_data), GLTL (GLTL-data), and Navi (Navi_data) datasets have been processed to include lifted natural language components by (NL2TL), and we perform similar processing of the Conformal dataset (conformalLTL_data) to include it in our evaluation.
 In Section 4.3 we develop and evaluate two grounding baselines on our three datasets.
@@
 First, we evaluate four language models on the natural language lifting task.
-The LLM-based approaches each use the lifting prompt template from the NL2TL framework [6], which includes few-shot ground-truth NL to lifted NL examples from each of the datasets.
-The input to both models is a natural-language sentence and we compare the prediction made by the model against the ground-truth lifted natural language and measure the average IoU scores over every individual NL AP.
-We present the results in Table 2 where we see the linguistic complexity of our datasets is highlighted in the accuracies as even the best scoring model (GPT-4.1) reduces in performance on our new datasets.
-While this performance drop is minor on this model, it is more significant on the lower-cost, smaller GPT models.
-Regardless, this indicates our success in increasing the evaluation complexity.
-Table 2: Comparison of lifting approaches. This table displays the mean IoU between predicted lifted natural language APs and ground-truth lifted natural language APs.
-Accuracy (%)
-S&R Traffic Light Warehouse
-Model GLTL [10] CW [18] Conformal [28] Navi [27] (ours) (ours) (ours)
+The LLM-based approaches each use the lifting prompt template from the NL2TL framework (NL2TL), which includes few-shot ground-truth NL to lifted NL examples from each of the datasets.
+The input to both models is a natural language sentence and we compare the prediction made by the model against the ground-truth lifted natural language using the lifting accuracy metric defined in Section 3.3.
+We present the results in Table 3 where we see the linguistic complexity of our datasets is highlighted in the accuracies, as even the best scoring model (GPT-4.1) reduces in performance on our new datasets.
+This performance drop is even more significant on the lower-cost, smaller GPT models.
+This indicates our success in increasing evaluation complexity.
+Table 3: Comparison of lifting approaches.
+Mean LiftAcc (%)
+S&R TL WH
+Model GLTL CW CF Navi (ours) (ours) (ours)
 GPT-3.5-turbo 81.6 78.6 76.8 71.0 65.3 59.4 67.9
@@
 4.2 Lifted Translation Evaluation
-Next, we evaluate the lifted translation capabilities of the three NL-to-LTL frameworks - nl2spec, NL2LTL, and NL2TL.
+Next, we evaluate the lifted translation capabilities of the three NL-to-LTL frameworks—nl2spec, NL2LTL, and NL2TL.
 In order to analyze the performance of their lifted translation abilities, the ground-truth lifted NL specification is given to the translation model, and the resulting lifted LTL translation is compared against the ground-truth lifted LTL.
-We perform this evaluation as a binary semantic accuracy using the pyModelChecking library [5] and present the results in Table 3.
+The formula for the translation accuracy metric is given in Section 3.3.
+We present these results in Table 4.
 Here, we see that lifted translation can be very successful with both out-of-the-box LLM prompting (nl2spec) and with fine-tuned seq2seq models.
 However, as we have noted, we will see in end-to-end evaluation that this is an overconfident estimation of translation performance as grounding is not considered.
-Table 3: Binary Accuracy on the Lifted NL to Lifted LTL task across seven datasets, for three frameworks (NL2LTL, nl2spec, NL2TL) given ground-truth lifted NL specifications.
-Accuracy (%)
-S&R Traffic Light Warehouse
-Framework Translation Model GLTL [10] CW [18] Conformal [28] Navi [27] (ours) (ours) (ours)
+Table 4: Comparison of four frameworks on the Lifted NL to Lifted TL translation task. Note that we provide ground-truth lifted NL specifications.
+TransAcc (%)
+S&R TL WH
+Framework Model GLTL CW CF Navi (ours) (ours) (ours)
 GPT-3.5-turbo 37.9 48.1 18.3 9.9 11.9 13.2 13.8
-NL2LTL [8] GPT-4o-mini 38.6 55.4 23.6 10.4 12.3 13.9 12.5
+NL2LTL (NL2LTL) GPT-4o-mini 38.6 55.4 23.6 10.4 12.3 13.9 12.5
 GPT-4.1-mini 51.7 64.6 42.1 39.7 41.6 40.0 37.4
 GPT-3.5-turbo 44.4 40.9 35.2 50.3 51.1 46.3 50.2
-nl2spec [7] GPT-4o-mini 77.3 80.1 73.5 69.7 74.9 75.8 74.2
+nl2spec (nl2spec) GPT-4o-mini 77.3 80.1 73.5 69.7 74.9 75.8 74.2
 GPT-4.1-mini 89.8 92.9 78.3 81.5 89.1 91.6 88.4
-NL2TL [6] t5-base 99.9 99.9 94.9 99.7 100.0 100.0 100.0
+NL2TL (NL2TL), Lang2LTL t5-base 99.9 99.9 94.9 99.7 100.0 100.0 100.0
 4.3 Grounding Evaluation
@@
 We use two prompting strategies (described in Appendix A.6) applied to three GPT models to provide a broad evaluation of current grounding capabilities.
-Our first prompting baseline - few-shot - is composed of a brief description of the task at hand, accompanied by nine few-shot examples of correct (sentence, lifted sentence, AP-dictionary) tuples from all three scenarios (as opposed to individual scenarios).
+Our first prompting baseline—few-shot—is composed of a brief description of the task at hand, accompanied by nine few-shot examples of correct (sentence, lifted sentence, AP-dictionary) tuples from all three scenarios (as opposed to individual scenarios).
 The next strategy is the scenario baseline prompt which includes the full scenario configuration file, as well as three few-shot examples from the dataset.
+Our final grounding baseline employs the same scenario-specific few-shot examples as the previous approach, with the addition of specific instructions to include intermediate reasoning steps used to arrive at the answer.
+We use GPT-4o for this chain-of-thought approach in order to evaluate the performance of reasoning-capable models on this task.
+All four models are instructed to format their final answer in JSON format.
 To measure grounding accuracy, we parse the resulting AP-dictionary predictions and compare them with our ground-truth knowledge of the AP-dictionary in each entry.
 Our metrics are per-AP and per-AP-dictionary accuracy.
-Per-AP accuracy is calculated by recording the total number of correctly grounded APs divided by the total number of APs in the test set, and per-AP-dictionary accuracy is calculated by recording the total number of correct complete AP-dictionaries, divided by the size of the test set.
+Per-AP accuracy is calculated by recording the total number of correctly grounded APs divided by the total number of APs in the test set, and per-AP-dictionary accuracy is calculated by recording the total number of completely correct AP-dictionaries, divided by the size of the test set.
 These results are presented in Table 5.
@@
 LLM Baseline uses 9 few–shot sentence + lifted sentence + AP dict examples from every dataset; “Scenario” includes the scenario definition in the prompt and 3 examples from only that dataset.
-Accuracy (of APs) Accuracy (of AP Dictionaries)
+Note that Lang2LTL grounds using cosine similarity between reference and canonical AP embeddings.
+Accuracy (% of APs) Accuracy (% of AP Dictionaries)
 Prompt Model S&R Traffic Light Warehouse S&R Traffic Light Warehouse
 GPT-3.5-turbo 56.9 69.5 18.3 34.2 51.4 7.4
-Few-shot GPT-4o-mini 82.3 66.5 18.4 68.6 48.4 7.0
+Few-shot General GPT-4o-mini 82.3 66.5 18.4 68.6 48.4 7.0
 GPT-4.1-mini 77.3 67.4 23.8 60.4 45.8 7.8
 GPT-3.5-turbo 76.7 37.3 13.6 63.6 20.8 5.0
-Scenario GPT-4o-mini 66.7 44.8 23.6 44.8 16.8 9.2
+Few-shot Scenario GPT-4o-mini 66.7 44.8 23.6 44.8 16.8 9.2
 GPT-4.1-mini 68.6 27.9 34.4 45.2 15.4 13.0
+Few-shot Chain-of-Thought GPT-4o 94.8 85.9 69.9 94.1 81.5 61.4
+Lang2LTL (Lang2LTL) N/A 77.6 86.2 61.8 59.0 73.6 38.8
 4.4 End-to-End Translation Evaluation
@@
 Framework S&R Traffic Light Warehouse
-NL2LTL [8] 35.4 38.4 26.2
-nl2spec [7] 34.8 33.6 29.6
-NL2TL [6] 54.4 60.1 46.2
+NL2LTL (NL2LTL) 35.4 38.4 26.2
+nl2spec (nl2spec) 34.8 33.6 29.6
+NL2TL (NL2TL) 54.4 60.1 46.2
+Lang2LTL (Lang2LTL) 58.5 72.1 37.9
 4.5 Verification Evaluation
@@
 Framework Grounding Strategy Sat Unsat Both Sat Unsat Both Sat Unsat Both
-NL2LTL [8] Few-shot 61.6 61.4 35.4 64.6 60.2 38.4 52.4 58.6 26.2
-Scenario 1.06 32.0 7.4 61.8 59.2 36.6 12.4 36.2 9.8
-nl2spec [7] Few-shot 47.4 48.0 34.8 47.2 46.0 33.6 46.0 44.2 29.6
-Scenario 34.0 36.4 21.0 40.2 41.8 28.2 32.0 34.6 19.0
-NL2TL [6] Few-shot 75.0 79.4 54.4 80.2 80.6 60.8 71.4 74.8 46.2
-Scenario 27.5 50.8 22.1 72.6 76.3 54.5 33.3 52.4 23.5
+NL2LTL (NL2LTL) Few-shot General 61.6 61.4 35.4 64.6 60.2 38.4 52.4 58.6 26.2
+Few-shot Scenario 1.06 32.0 7.4 61.8 59.2 36.6 12.4 36.2 9.8
+nl2spec (nl2spec) Few-shot General 47.4 48.0 34.8 47.2 46.0 33.6 46.0 44.2 29.6
+Few-shot Scenario 34.0 36.4 21.0 40.2 41.8 28.2 32.0 34.6 19.0
+NL2TL (NL2TL) Few-shot General 75.0 79.4 54.4 80.2 80.6 60.8 71.4 74.8 46.2
+Few-shot Scenario 27.5 50.8 22.1 72.6 76.3 54.5 33.3 52.4 23.5
+Lang2LTL (Lang2LTL) Embedding 43.3 61.9 39.3 44.7 63.0 41.3 21.6 40.1 16.6
 5 Conclusion
@@
 We acknowledge that the datasets included in the VLTL-Bench suite are generated using a finite number of linguistic and logical templates, populated by diverse synthetic natural language APs.
-Therefore, our logical diversity is limited to 43 lifted NL-LTL templates.
-The linguistic complexity of our benchmark is also partially limited in the same manner.
-However, our randomly selected action and target references combined with LLM-based pruning allows this work to surpass existing datasets in terms of diversity and realism, and we believe future effort should be partially devoted to further improvements in this area.
 VLTL-Bench reveals significant weaknesses in what were previously ironclad NL-to-LTL translation frameworks.
-Among these our the reliance on accurately lifted NL inputs for translation, lack of accurate grounding components, and lack of example trace inputs in current approaches.
-We hope that contribution will encourage exploration of diverse methods for NL-to-LTL translation, beyond the use of LLMs.
+Among these weakness are: the reliance on accurately lifted NL inputs for translation, lack of accurate grounding components, and lack of example trace inputs in current approaches.
+We envision our contribution will encourage exploration of diverse methods for grounded NL-to-LTL translation, beyond the use of LLMs.
 Appendix A Appendix
-In this appendix, we present a detailed overview of linear temporal logic in A.1, a discussion of verification via Kripke structures in A.2, a quantitative comparison of our VLTL-Bench dataset against existing datasets as well as examples from those datasets in A.3, our developed prompts for the baseline grounding approaches in A.4, the configuration files for our three scenarios in A.5, and finally our estimated compute resource usage and our external code and license information in A.6.
+In this appendix, we present a detailed overview of linear temporal logic in A.1, a discussion of verification via Kripke structures in A.2, a quantitative comparison of our VLTL-Bench dataset against existing datasets as well as examples from those datasets in A.5, our developed prompts for the baseline grounding approaches in A.6, the configuration files for our three scenarios in A.8, and finally our estimated compute resource usage and our external code and license information in A.9.
 A.1 Linear Temporal Logic
-Linear temporal logic (LTL) is a modal extension of classical propositional logic that enables reasoning about how truths evolve over a discrete, linear timeline [32].
+Linear temporal logic (LTL) is a modal extension of classical propositional logic that enables reasoning about how truths evolve over a discrete, linear timeline (LTL).
 Formulas in LTL are interpreted over infinite sequences (or “traces”) of states
@@
 $\sigma,i\models\varphi_{1}\cup\varphi_{2}$ $\quad\text{iff}\quad\exists k\geq i:\;\sigma,k\models\varphi_{2}\;\wedge\;\forall j\in[i,k):\;\sigma,j\models\varphi_{1}.$
-Although our focus is on discrete-time LTL, many of these ideas carry over to related formalisms such as signal temporal logic (STL) for continuous-time, real-valued signals [19].
+Although our focus is on discrete-time LTL, many of these ideas carry over to related formalisms such as signal temporal logic (STL) for continuous-time, real-valued signals (STL).
 A.2 Verification via Kripke Structures and Fluents
 Verification of LTL specifications is typically conducted using a Kripke structure, which is a formal transition system comprising states, transitions, and labels indicating which atomic propositions hold true in each state.
-Formally, a Kripke structure is defined as a tuple $M=(S,S\_0,R,L)$, where:
+Formally, a Kripke structure is defined as a tuple M=(S,S_0,R,L)M=(S,S\_0,R,L), where:
   • $S$ is a finite set of states,
@@
   • $R\subseteq S\times S$ is the transition relation, specifying allowed state transitions,
-[…]
+  • $L:S\rightarrow 2^{\mathit{AP}}$ is a labeling function mapping states to the sets of atomic propositions that are true in each state.
 Verification involves checking whether every possible path through the Kripke structure satisfies the given LTL formula.
@@
 Fluents bridge the gap between high-level temporal specifications and lower-level state transitions, facilitating practical model checking and control synthesis in robot control systems.
-A.3 Existing Datasets
-Table 7: Comparison of NL–LTL datasets.
-Note that VLTL-Bench datasets include satisfying and unsatisfying traces for each entry.
-VLTL-Bench separates the total data into training and testing, as well as a subset of data taken from training and testing partitions that was selected by GPT-4o for its particularly realistic natural language.
-Dataset Size Unique TL # APs # Actions # Args
-GLTL Gopalan et al. [10] 11,109 37 4 1 4
-CW MacGlashan et al. [18] 3,371 37 4 1 4
-Conformal Wang et al. [28] 1,000 212 239 4 235
-Navi Wang et al. [27] 7,474 6,414 221 26
-Search-and-rescue [VLTL-Bench] 7,304 5,425 220 7 44
-Traffic-light [VLTL-Bench] 7,319 6,196 5,046 4 175
-Warehouse [VLTL-Bench] 7,457 5,991 5,074 5 452
+A.3 VLTL-Bench LTL Expression Statistics
+Token / Operator Search & Rescue Traffic Light Warehouse
+and 5536 5508 5297
+double_implies 1104 1141 1091
+finally 3835 3842 3918
+globally 9899 9851 9820
+implies 5164 5295 5264
+next 12144 12304 11713
+not 6781 6724 6593
+or 3198 3229 3054
+prop_1 14440 14466 14193
+prop_2 7934 7964 7835
+prop_3 3740 3831 3825
+until 1112 1088 1147
+Table 8: Operator splits and template breakdowns by domain.
+A.4 VLTL-Bench New Templates
+We then craft 7 of our own templates to fill perceived gaps in specification coverage.
+Of these templates, 4 entries include new lifted LTL halves (marked below with a *), and 3 include new lifted NL halves.
+NL LTL
+finally ( not prop_1) “eventually, avoid prop_1”
+globally ( not prop_1) “always avoid prop_1”; “prop_1 must never occur”
+next prop_1 “at the next time step, prop_1 holds”
+prop_1 until prop_2 “prop_1 must always hold at all times before prop_2”
+finally (prop_1 and prop_2) OLD: “Eventually, both prop_1 and prop_2 will hold simultaneously”
+NEW: “At some point, prop_1 and prop_2 will both hold at the same time.”
+globally (prop_1 and prop_2) OLD: “Both prop_1 and prop_2 hold at every step.”
+NEW: “At all time steps, prop_1 and prop_2 both hold.”
+finally (prop_1 or prop_2) OLD: “eventually, either prop_1 or prop_2”
+NEW: “either prop_1 or prop_2 will hold at some point in time.”
+Table 9: Examples of NL–LTL mappings. OLD/NEW entries show updated phrasing.
+A.5 Existing Datasets
 Cleanup World (CW).
@@
   • APs: prop_1 = go to parking lot 4, prop_2 = go to car 5
-A.4 Grounding Prompts
+A.6 Grounding Prompts
 This section includes the few-shot examples used in our grounding prompt baselines.
 The few-shot baselines uses all of the following in its prompt, while the scenario baseline includes only the scenario specific few-shot examples combined with the scenario description, given in Appendix A.8
-Figure 3: Few-shot examples for the Warehouse Scenario
-Figure 4: Few-shot examples for the Traffic Light Scenario
-Figure 5: Few-shot examples for the Search and Rescue Scenario
-A.5 Scenario Configurations
+Few-shot Prompt:
+”role”: ”system”, ”content”: ”You are an LTL translation assistant, your goal is to return the desired prop_dict, a dictionary that relates natural language atomic proposition/predicate references to their canonical/known representation in the scenario.”, ”role”: ”user”, ”content”: Few-shot Examples: {examples from ALL domains, shown in appendix A.7, total of 9 examples} Now predict: Sentence: {sentence} Lifted: {lifted_sentence} Prop_dict:
+Scenario Prompt:
+”role”: ”system”, ”content”: ”You are an LTL translation assistant, your goal is to return the desired prop_dict, a dictionary that relates natural language atomic proposition/predicate references to their canonical/known representation in the scenario.”, ”role”: ”user”, ”content”: Scenario Configuration: scenario yaml, given in appendix A.8 Few-shot Examples: {examples from this specific scenario, shown in Appendix A.7} Now predict: Sentence: {sentence} Lifted: {lifted_sentence} Prop_dict:
+A.7 Few-shot Examples by Scenario
+Warehouse Examples
+Sentence: [”The system must eventually, avoid prop_1”] Lifted Sentence: [”The system must eventually, avoid prop_1”] prop_dict: { ”prop_1”: { ”action_canon”: ”deliver”, ”action_ref”: ”drop off”, ”args_canon”: [”sandwich loading_dock”], ”args_ref”: [”square food loading dock”] } } Sentence: [”Whenever prop_1 holds, prop_2 holds as well.”] Lifted Sentence: [”Whenever prop_1 holds, prop_2 holds as well.”] prop_dict: { ”prop_1”: { ”action_canon”: ”idle”, ”action_ref”: ”remain still”, ”args_canon”: [], ”args_ref”: [] }, ”prop_2”: { ”action_canon”: ”get_help”, ”action_ref”: ”call for help”, ”args_canon”: [], ”args_ref”: [] } } Sentence: [”If prop_2 holds, then in the next step prop_3 persists until prop_1 holds, or else prop_3 holds forever.”] Lifted Sentence: [”If prop_2 holds, then in the next step prop_3 persists until prop_1 holds, or else prop_3 holds forever.”] prop_dict: { ”prop_1”: { ”action_canon”: ”pickup”, ”action_ref”: ”grab”, ”args_canon”: [”hot_dog”], ”args_ref”: [”bunned sausage”] }, ”prop_2”: { ”action_canon”: ”pickup”, ”action_ref”: ”grab”, ”args_canon”: [”potted_plant”], ”args_ref”: [”plant”] }, ”prop_3”: { ”action_canon”: ”search”, ”action_ref”: ”search for”, ”args_canon”: [”cup”], ”args_ref”: [”beverage cup”] } } Search and Rescue Examples
+Sentence: [”This controller must always avoid prop_1”] Lifted Sentence: [”This controller must always avoid prop_1”] prop_dict: { ”prop_1”: { ”action_canon”: ”record”, ”action_ref”: ”begin recording”, ”args_canon”: [”fire_source”], ”args_ref”: [”fire source”] } } Sentence: [”In this task, take a photo of flood, then return home.”] Lifted Sentence: [”In this task, prop_1 then prop_2”] prop_dict: { ”prop_1”: { ”action_canon”: ”photo”, ”action_ref”: ”take a photo of”, ”args_canon”: [”flood”], ”args_ref”: [”flood”] }, ”prop_2”: { ”action_canon”: ”go_home”, ”action_ref”: ”return home”, ”args_canon”: [], ”args_ref”: [] } } Sentence: [”If every record flood is eventually followed by talking to the safe victim, then avoid the impending debris must occur infinitely often.”] Lifted Sentence: [”If every prop_1 is eventually followed by prop_2 then prop_3 must occur infinitely often.”] prop_dict: { ”prop_1”: { ”action_canon”: ”record”, ”action_ref”: ”record”, ”args_canon”: [”flood”], ”args_ref”: [”flood”] }, ”prop_2”: { ”action_canon”: ”communicate”, ”action_ref”: ”talk to”, ”args_canon”: [”safe_victim”], ”args_ref”: [”safe victim”] }, ”prop_3”: { ”action_canon”: ”avoid”, ”action_ref”: ”avoid”, ”args_canon”: [”impending_debris”], ”args_ref”: [”impending debris”] } }
+Traffic Light Examples
+Sentence: [”You”, ”must”, ”eventually,”, ”avoid”, ”set”, ”east”, ”light”, ”yellow.”] Grounded: [”You”, ”must”, ”eventually,”, ”avoid”, ”prop_1”] prop_dict: { ”prop_1”: { ”action_canon”: ”change”, ”action_ref”: ”set”, ”args_canon”: [”light_east”, ”yellow”], ”args_ref”: [”east light”, ”yellow”] } } Sentence: [”Both”, ”change”, ”west”, ”light”, ”red”, ”and”, ”take”, ”a”, ”video”, ”of”, ”the”, ”car”, ”on”, ”southwest”, ”10th”, ”avenue”, ”hold”, ”at”, ”every”, ”step.”] Grounded: [”Both”, ”prop_1”, ”and”, ”prop_2”, ”hold”, ”at”, ”every”, ”step.”] prop_dict: { ”prop_1”: { ”action_canon”: ”change”, ”action_ref”: ”change”, ”args_canon”: [”light_west”, ”red”], ”args_ref”: [”west light”, ”red”] }, ”prop_2”: { ”action_canon”: ”record”, ”action_ref”: ”take a video of”, ”args_canon”: [”car”, ”southwest_10th_avenue”], ”args_ref”: [”car”, ”southwest 10th avenue”] } } Sentence: [”If”, ”take”, ”a”, ”picture”, ”of”, ”the”, ”car”, ”on”, ”northwest”, ”6th”, ”street”, ”holds”, ”and”, ”set”, ”east”, ”light”, ”green”, ”holds”, ”next,”, ”then”, ”request”, ”assistance”, ”holds”, ”in”, ”the”, ”step”, ”after”, ”that.”] Grounded: [”If”, ”prop_1”, ”holds”, ”and”, ”prop_2”, ”holds”, ”next,”, ”then”, ”prop_3”, ”holds”, ”in”, ”the”, ”step”, ”after”, ”that.”] prop_dict: { ”prop_1”: { ”action_canon”: ”photo”, ”action_ref”: ”take a picture of”, ”args_canon”: [”car”, ”northwest_6th_street”], ”args_ref”: [”car”, ”northwest 6th street”] }, ”prop_2”: { ”action_canon”: ”change”, ”action_ref”: ”set”, ”args_canon”: [”light_east”, ”green”], ”args_ref”: [”east light”, ”green”] }, ”prop_3”: { ”action_canon”: ”get_help”, ”action_ref”: ”request assistance”, ”args_canon”: [], ”args_ref”: [] } }
+A.8 Scenario Configurations
 In this section, we provide the scenario configuration files that are inserted into the grounding prompts and used for data generation.
@@
 Figure 5: Search and Rescue Scenario Configuration file
-A.6 Compute Resources and External Code and License Information
+Figure 6: Kitchen Assistant Scenario Configuration file
+A.9 Compute Resources and External Code and License Information
 All LLM inference was performed using the OpenAI API.
@@
 The nl2spec framework is released at https://github.com/realChrisHahn2/nl2spec under the MIT license, the NL2TL framework is released at https://github.com/yongchao98/NL2TL?tab=readme-ov-file with no attached license, the NL2LTL framework is released at https://github.com/IBM/nl2ltl under the MIT license, and the pyModelChecking library is released at https://github.com/albertocasagrande/pyModelChecking under the GNU General Public License.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: All claims in the beginning of the paper are justified by the experiments and details provided in the main body of the paper.
-  5. 2.
-    Limitations
-  6. Question: Does the paper discuss the limitations of the work performed by the authors?
-  7. Answer: [Yes]
-  8. Justification: Limitations are discussed in Section 5.
-  9. 3.
-    Theory assumptions and proofs
-  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  11. Answer: [N/A]
-  12. Justification: We do not provide any theories. All results are experimental, and all experiments are provided.
-  13. 4.
-    Experimental result reproducibility
-  14.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  15. Answer: [Yes]
-  16.
-Justification: The paper is thorough in its description of the experiments and the data.
-Someone familiar with the area could feasibly replicate this work given the citations to existing work and the details provided.
-Regardless, we provide all code and the dataset is public and provided as required.
-  17. 5.
-    Open access to data and code
-  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  19. Answer: [Yes]
-  20. Justification: The dataset is public and the code and experimental details will be in the supplemental material.
-  21. 6.
-    Experimental setting/details
-  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  23. Answer: [Yes]
-  24. Justification: Full details are in the code, but the information in the paper is sufficient for understanding the work and the meaning of the evaluations.
-  25. 7.
-    Experiment statistical significance
-  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  27. Answer: [No]
-  28.
-Justification: Our evaluations present only the mean of each metric.
-Our evaluations are (relatively) small, composed of <1000 examples, and we focus on the presentation and descriptions of our new datasets, rather than a detailed review of current frameworks.
-  29. 8.
-    Experiments compute resources
-  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  31. Answer: [Yes]
-  32. Justification: See Appendix A.6.
-  33. 9.
-    Code of ethics
-  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  35. Answer: [Yes]
-  36. Justification: We have no ethical issues and have read the code.
-  37. 10.
-    Broader impacts
-  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  39. Answer: [Yes]
-  40. Justification: We discuss the benefits to NL to TL translation which can be used for improved safe operation of formally specified systems.
-  41. 11.
-    Safeguards
-  42.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  43. Answer: [N/A] .
-  44. Justification: Our dataset does not enable dangerous models or model behavior.
-  45. 12.
-    Licenses for existing assets
-  46.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  47. Answer: [Yes]
-  48. Justification: See Appendix A.6.
-  49. 13.
-    New assets
-  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  51. Answer: [Yes] .
-  52. Justification: Dataset is published as required. All code released in the supplemental material is well documented.
-  53. 14.
-    Crowdsourcing and research with human subjects
-  54.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  55. Answer: [N/A]
-  56. Justification: No human subjects
-  57. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  58.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  59. Answer: [N/A]
-  60. Justification: No human subjects
-  61. 16.
-    Declaration of LLM usage
-  62.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  63. Answer: [Yes]
-  64.
-Justification: All uses of LLM for help with dataset generation and for experimental evaluation of methods which use LLMs were disclosed.
-In addition we have provided the prompts for our unique usage.
+A.10 Large Language Model Disclosure
+During the preparation of this paper, the authors employed large language models (LLMs) as assistive tools for limited tasks including proof-reading, text summarization, and the discovery of related work.
+All substantive research contributions, analyses, and claims presented in this paper were conceived, developed, and verified by the authors.
+The authors maintain full ownership and responsibility for the content of the paper, including its technical correctness, originality, and scholarly contributions.